### v1

# Erdélyi-type integrals for $F_{K}$ function and their $q$-analogues

Abstract
In this paper, we revisit the recent result of Luo, Xu, and Raina [Fractal Frac. 6 (3) (2022)] on an Erdélyi-type integral for Saran’s three-variable hypergeometric function $F_{K}$. We provide a new proof of this integral and derive an attractive new integral related to Appell’s function $F_{2}$. A further extension on the $L$-variable $F_{K}$ function, which appears in physics, is also discussed. Furthermore, we prove various $q$-Erdélyi-type integrals for the $q$-analogue of the $F_{K}$-function. An interesting discrete analogue is also included. We also provide a valuable compilation of the sources for known Erdélyi-type integrals of many different hypergeometric functions in the Appendix.
Keywords: Erdélyi-type integral, fractional integration by parts, hypergeometric function, $q$-analogue, Saran’s function.
Mathematics Subject Classification (2020): 33C05, 33C65, 33C70, 33D15, 33D70.
1School of Mathematics and Statistics,
Donghua University, Shanghai 201620,
People’s Republic of China.
E-mail: mathwinnie@live.com, mathwinnie@dhu.edu.cn
2School of Mathematics and Statistics,
Donghua University, Shanghai 201620,
People’s Republic of China.
E-mail: guoliangjia77@outlook.com
1 Introduction
1.1 Background and motivation
The Gauss hypergeometric function $_{2}F_{1}$ is defined by
$_{2}F_{1}\left[\begin{matrix}\alpha,\beta\\ \gamma\end{matrix};z\right]:=\sum_{n=0}^{\infty}\frac{(\alpha)_{n}(\beta)_{n}}{(\gamma)_{n}}\frac{z^{n}}{n!}\penalty 10000\ \penalty 10000\ \penalty 10000\ (|z|<1),$
where $\alpha,\beta\in\mathbb{C}$ and $\gamma\notin\mathbb{Z}_{\leq 0}:=\{0,-1,-2,\cdots\}$ (see [13, p. 56] and [38, p. 384]). It was shown by Euler that [13, p. 59, Eq. (10)]
${}_{2}F_{1}\left[\begin{matrix}\alpha,\beta\\ \gamma\end{matrix};z\right]=$ $\int_{0}^{1}(1-zt)^{-\alpha}\mathrm{d}\mu_{\beta,\gamma-\beta}(t)\penalty 10000\ \penalty 10000\ \penalty 10000\ (\Re(\gamma)>\Re(\beta)>0),$ (1)
${}_{2}F_{1}\left[\begin{matrix}\alpha,\beta\\ \gamma\end{matrix};z\right]=$ $\int_{0}^{1}(1-zt)^{-\beta}\mathrm{d}\mu_{\alpha,\gamma-\alpha}(t)\penalty 10000\ \penalty 10000\ \penalty 10000\ (\Re(\gamma)>\Re(\alpha)>0).$ (2)
Here and throughout the paper, we let $\mu_{\alpha,\beta}$ be the Dirichlet measure defined by [6, p. 64, Definition 4.4-1]
$\mathrm{d}\mu_{\alpha,\beta}(t):=\mathfrak{m}_{\alpha,\beta}(t)\mathrm{d}t,$ (3)
where
$\mathfrak{m}_{\alpha,\beta}(t):=\frac{\Gamma(\alpha+\beta)}{\Gamma(\alpha)\Gamma(\beta)}t^{\alpha-1}\left(1-t\right)^{\beta-1}\penalty 10000\ \penalty 10000\ \penalty 10000\ (\min\{\Re(\alpha),\Re(\beta)\}>0).$
Obviously, $\int_{0}^{1}\mathrm{d}\mu_{\alpha,\beta}(t)=1$. In general, integrals (1) and (2) cannot be directly transformed into each other, but Riemann [41] indicated that both could be derived from a multiple integral. According to the existing literature, Schellenberg [46], Wirtinger [52], Poole [5] and Erdélyi [11] studied this problem. We are particularly interested in Erdélyi’s and Poole’s methods.
Erdélyi’s method (see [11, p. 272–273]) is to substitute
$(1-zt)^{-\alpha}=\int_{0}^{1}(1-zts)^{-\gamma}\mathrm{d}\mu_{\alpha,\gamma-\alpha}(s)\penalty 10000\ \penalty 10000\ \penalty 10000\ (\Re(\gamma)>\Re(\alpha)>0)$
into the first integral (1) to obtain the double integral
$\int_{0}^{1}\int_{0}^{1}(1-zts)^{-\gamma}\mathrm{d}\mu_{\beta,\gamma-\beta}(t)\mathrm{d}\mu_{\alpha,\gamma-\alpha}(s),$
which is symmetric in $\alpha$ and $\beta$. If we integrate with respect to $t$ first, then we obtain the second integral (2).
Poole’s method [5] is to apply the integration by parts to the contour integral representation of $_{2}F_{1}$, which is very delightful. After reading Poole’s paper, Erdélyi [10] realized that Poole’s method can be generalized by using the fractional derivatives. By inventing the method of fractional integration by parts, Erdélyi not only gave a simple description of Poole’s derivation but also reproduced the Bateman integral
$_{2}F_{1}\left[\begin{matrix}\alpha,\beta\\ \gamma\end{matrix};z\right]=\int_{0}^{1}{}_{2}F_{1}\left[\begin{matrix}\alpha,\beta\\ \lambda\end{matrix};zx\right]\mathrm{d}\mu_{\lambda,\gamma-\lambda}(x)\hskip 14.22636pt(\Re(\gamma)>\Re(\lambda)>0).$ (4)
In [12], Erdélyi further developed his method and proved the following important integrals for $_{2}F_{1}$:
${}_{2}F_{1}\left[\begin{matrix}\alpha,\beta\\ \gamma\end{matrix};z\right]$ $=\int_{0}^{1}\left(1-zx\right)^{-\alpha^{\prime}}{}_{2}F_{1}\left[\begin{matrix}\alpha-\alpha^{\prime},\beta\\ \lambda\end{matrix};zx\right]$
$\hskip 14.22636pt\cdot{}_{2}F_{1}\left[\begin{matrix}\alpha^{\prime},\beta-\lambda\\ \gamma-\lambda\end{matrix};\frac{(1-x)z}{1-xz}\right]\mathrm{d}\mu_{\lambda,\gamma-\lambda}(x)\hskip 14.22636pt(\Re(\gamma)>\Re(\lambda)>0),$ (5)
${}_{2}F_{1}\left[\begin{matrix}\alpha,\beta\\ \gamma\end{matrix};z\right]$ $=\int_{0}^{1}\left(1-zx\right)^{\lambda-\alpha-\beta}{}_{2}F_{1}\left[\begin{matrix}\lambda-\alpha,\lambda-\beta\\ \eta\end{matrix};zx\right]$
$\hskip 14.22636pt\cdot{}_{2}F_{1}\left[\begin{matrix}\alpha+\beta-\lambda,\lambda-\eta\\ \gamma-\eta\end{matrix};\frac{(1-x)z}{1-xz}\right]\mathrm{d}\mu_{\eta,\gamma-\eta}(x)\hskip 14.22636pt(\Re(\gamma)>\Re(\eta)>0),$ (6)
${}_{2}F_{1}\left[\begin{matrix}\alpha,\beta\\ \gamma\end{matrix};z\right]$ $=\int_{0}^{1}{}_{3}F_{2}\left[\begin{matrix}\alpha,\beta,\eta\\ \lambda,\nu\end{matrix};zx\right]\mathrm{d}\mu_{\eta-\lambda,\gamma-\lambda,\gamma-\lambda+\eta-\nu,\nu}(x)$
$\hskip 28.45274pt(\min\{\Re(\nu),\Re(\lambda),\Re(\gamma-\lambda+\eta-\nu)\}>0).$ (7)
In (7), $\mu_{\alpha,\beta,\gamma,\eta}$ is the hypergeometric measure defined by [32, p. 6, Eq. (11)]
$\mathrm{d}\mu_{\alpha,\beta,\gamma,\eta}(t):=\mathfrak{m}_{\alpha,\beta,\gamma,\eta}(t)\mathrm{d}t,$ (8)
where $\min\{\Re(\eta),\Re(\gamma),\Re(\eta+\gamma-\alpha-\beta)\}>0$ and
$\mathfrak{m}_{\alpha,\beta,\gamma,\eta}(t)$ $:=\frac{\Gamma(\eta+\gamma-\alpha)\Gamma(\eta+\gamma-\beta)}{\Gamma(\eta)\Gamma(\gamma)\Gamma(\eta+\gamma-\alpha-\beta)}t^{\eta-1}(1-t)^{\gamma-1}{}_{2}F_{1}\left[\begin{matrix}\alpha,\beta\\ \gamma\end{matrix};1-t\right].$
We have $\int_{0}^{1}\mathrm{d}\mu_{\alpha,\beta,\gamma,\eta}(t)=1$ and $\mathfrak{m}_{0,\beta,\gamma,\eta}(t)=\mathfrak{m}_{\alpha,0,\gamma,\eta}(t)=\mathfrak{m}_{\eta,\gamma}(t)$.
It is particularly noteworthy that Erdélyi’s integrals are not only of computational interest but also have many important and profound applications. Gasper [17] has shown that (6) provides a natural way to the formulas of Dirichlet-Mehler type for the Jacobi polynomials and the generalized Legendre functions. Sprinkhuizen-Kuyper [47] used (6) to prove a compositional property for her fractional integral operator $I_{\nu}^{\mu,\lambda}$. The integral (6) also plays an important role in the proof of product formula for biangle polynomials [25]. The proofs of the compositional properties of Saigo’s fractional integral operators $I^{\alpha,\beta,\eta}$ and $J^{\alpha,\beta,\eta}$ are based on integrals (5) and (6) [44]. Raina [39] showed that the integral (5) can be used to construct a solution to a certain Abel-type integral equation involving the Appell hypergeometric function $F_{3}$ in the kernel.
We refer to integrals that have a form similar to integrals (5)–(7) as Erdélyi-type integrals. More precisely, for a given hypergeometric function, its Erdélyi-type integrals can generally express it as an integral of the product of two functions of the same type (see [32, p. 2]).
Many Erdélyi-type integrals for various kinds of univariate and multivariate hypergeometric functions have been discovered using the technique of Erdélyi’s fractional integration by parts. For example, Luo and Raina [30] recenlty extended (6) to a special class of generalized hypergeometric functions by using the theory of Miller and Paris and then found their applications in the theory of certain generalized fractional integral operator [33]. The table in Appendix 5 lists all hypergeometric functions for which Erdélyi-type integrals and some of their $q$- and discrete analogues have been established.
1.2 The plan of the work
Very recently, there has been a sustained growth in research interest regarding the three-variable hypergeometric function $F_{K}$ introduced by Saran [45] and many interesting results have been achieved (see [1], [8], [20], [31] and [32]). The $F_{K}$-function is defined by
$F_{K}[\alpha_{1},\alpha_{2},\alpha_{2},\beta_{1},\beta_{2},\beta_{1};\gamma_{1},\gamma_{2},\gamma_{3};x,y,z]$
$\hskip 14.22636pt:=\sum_{m,n,p=0}^{\infty}\frac{(\alpha_{1})_{m}(\alpha_{2})_{n+p}(\beta_{1})_{m+p}(\beta_{2})_{n}}{(\gamma_{1})_{m}(\gamma_{2})_{n}(\gamma_{3})_{p}m!n!p!}x^{m}y^{n}z^{p}$ (9)
$\hskip 14.22636pt=\sum_{p=0}^{\infty}\frac{(\alpha_{2})_{p}(\beta_{1})_{p}}{p!(\gamma_{3})_{p}}{}_{2}F_{1}\left[\begin{matrix}\beta_{1}+p,\alpha_{1}\\ \gamma_{1}\end{matrix};x\right]{}_{2}F_{1}\left[\begin{matrix}\alpha_{2}+p,\beta_{2}\\ \gamma_{2}\end{matrix};y\right]z^{p},$ (10)
where $(x,y,z)\in\mathbb{D}_{K}:=\{(x,y,z)\in\mathbb{C}^{3}:|x|<1,|y|<1,|z|<(1-|x|)(1-|y|)\}$.
In 2022, Luo et al. [32] discovered an elegant Erdélyi-type integral for Saran’s $F_{K}$-function.
Theorem 1.1 ([32]).
Let $\Re(\alpha_{1}+\eta_{1})>\Re(\lambda_{1})>0$, $\Re(\beta_{2}+\mu_{2})>\Re(\lambda_{2})>0$ and $\Re(\gamma_{3})>\Re(\beta_{1})>0$. Then we have
$F_{K}[\alpha_{1},\alpha_{2},\alpha_{2},\beta_{1},\beta_{2},\beta_{1};\alpha_{1}+\eta_{1},\beta_{2}+\mu_{2},\gamma_{3};x,y,z]=\int_{0}^{1}\int_{0}^{1}\int_{0}^{1}(1-ux)^{-\lambda_{3}}(1-vy)^{-\eta_{2}}$
$\hskip 14.22636pt\cdot F_{K}[\alpha_{1},\alpha_{2}-\eta_{2},\alpha_{2}-\eta_{2},\beta_{1}-\lambda_{3},\beta_{2},\beta_{1}-\lambda_{3};\alpha_{1}-\lambda_{1}+\eta_{1},\beta_{2}-\lambda_{2}+\mu_{2},\beta_{1}-\lambda_{3};ux,vy,wz]$
$\hskip 14.22636pt\cdot F_{K}\left[\lambda_{1}-\eta_{1},\eta_{2},\eta_{2},\lambda_{3},\lambda_{2}-\mu_{2},\lambda_{3};\lambda_{1},\lambda_{2},\lambda_{3};\frac{(1-u)x}{1-ux},\frac{(1-v)y}{1-vy},\frac{wz}{(1-ux)(1-vy)}\right]$
$\hskip 14.22636pt\cdot\mathrm{d}\mu_{\alpha_{1}-\lambda_{1}+\eta_{1},\lambda_{1}}(u)\mathrm{d}\mu_{\beta_{2}-\lambda_{2}+\mu_{2},\lambda_{2}}(v)\mathrm{d}\mu_{\beta_{1},\gamma_{3}-\beta_{1}}(w),$ (11)
where $(x,y,z)\in\mathbb{D}_{K}$ and $\mathrm{d}\mu_{\alpha,\beta}(t)$ is defined in (3).
However, the initial proof of integral (1.1) was relatively difficult. In fact, Luo et al. first established a very general integral involving the Srivastava-Daoust function through fractional integration by parts, and then reduced it to (1.1) by specializing the parameters. In this paper, we shall provide a direct and simple proof of Theorem 1.1 in Section 3.1. Then in Section 3.2, we shall derive a curious integral representation for Appell function $F_{2}$ from Theorem 1.1. A further extension, which is related to Belitsky’s generalization of $F_{K}$ [4, p. 69], is considered in Section 3.3.
Section 4 devotes to the various $q$-Erdélyi-type integrals for $q$-analogue of $F_{K}$. Throughout this paper, we assume that $0<q<1$. For $a\in\mathbb{C}$, we define the $q$-shifted factorials as follows
$(a;q)_{0}=1,\penalty 10000\ \penalty 10000\ \penalty 10000\ (a;q)_{n}=\prod_{j=0}^{n-1}(1-aq^{j})\penalty 10000\ \penalty 10000\ \penalty 10000\ (n\in\mathbb{Z}_{\geq 1})$
and
$(a;q)_{\infty}=\lim_{n\rightarrow+\infty}(a;q)_{n}=\prod_{j=0}^{\infty}(1-aq^{j}).$
The multiple $q$-shifted factorials are defined by $(a_{1},\cdots,a_{m};q)_{n}=(a_{1};q)_{n}\cdots(a_{m};q)_{n}$, where $n$ could be an integer or $\infty$. Let $f:\mathbb{C}^{k}\rightarrow\mathbb{C}$ be a function. The $k$-dimensional $q$-integral can be defined by (see, for example, [23, p. 1] and [51, p. 241])
$\int_{[0,1]^{k}}f(\mathbf{x})\mathrm{d}_{q}\mathbf{x}=(1-q)^{k}\sum_{\mathbf{n}\in\mathbb{Z}_{\geq 0}^{k}}f(q^{\mathbf{n}})q^{\langle\mathbf{n}\rangle}.$ (12)
where $\mathbf{x}=(x_{1},\cdots,x_{k})$, $\sum_{\mathbf{n}\in\mathbb{Z}_{\geq 0}^{k}}\equiv\sum_{n_{1}=0}^{\infty}\cdots\sum_{n_{k}=0}^{\infty}$ and $\langle\mathbf{n}\rangle=n_{1}+\cdots+n_{k}$.
The $q$-gamma function $\Gamma_{q}(x)$ is defined by [19, p. 20, Eq. (1.10.1)]
$\Gamma_{q}(x)=\frac{(q;q)_{\infty}}{(q^{x};q)_{\infty}}(1-q)^{1-x}$
and the $q$-beta function is defined by [19, p. 22, Eq. (1.10.13)]
$B_{q}(x,y)=\frac{\Gamma_{q}(x)\Gamma_{q}(y)}{\Gamma_{q}(x+y)}=(1-q)\frac{(q,q^{x+y};q)_{\infty}}{(q^{x},q^{y};q)_{\infty}}.$
Note that $\lim_{q\rightarrow 1^{-}}\Gamma_{q}(x)=\Gamma(x)$ and $\lim_{q\rightarrow 1^{-}}B_{q}(x,y)=B(x,y)$. The $q$-integral representation for $B_{q}(x,y)$ is given by [19, p. 23, Eq. (1.11.7)]
$B_{q}(x,y)=\int_{0}^{1}t^{x-1}\frac{(tq;q)_{\infty}}{(tq^{y};q)_{\infty}}\mathrm{d}_{q}t,\penalty 10000\ \penalty 10000\ \penalty 10000\ \Re(x)>0,\penalty 10000\ y\notin\mathbb{Z}_{\leq 0}.$
The $q$-analogue of $F_{K}$ can be defined by (see [15] and [43])
$\Phi_{K}[\alpha_{1},\alpha_{2},\alpha_{2},\beta_{1},\beta_{2},\beta_{1};\gamma_{1},\gamma_{2},\gamma_{3};q,x,y,z]$
$\hskip 14.22636pt:=\sum_{m,n,p=0}^{\infty}\frac{(\alpha_{1};q)_{m}(\alpha_{2};q)_{n+p}(\beta_{1};q)_{m+p}(\beta_{2};q)_{n}}{(\gamma_{1},q;q)_{m}(\gamma_{2},q;q)_{n}(\gamma_{3},q;q)_{p}}x^{m}y^{n}z^{p}$ (13)
$\hskip 14.22636pt=\sum_{p=0}^{\infty}\frac{(\alpha_{2},\beta_{1};q)_{p}}{(\gamma_{3},q;q)_{p}}{}_{2}\phi_{1}\left[\begin{matrix}\beta_{1}q^{p},\alpha_{1}\\ \gamma_{1}\end{matrix};q,x\right]{}_{2}\phi_{1}\left[\begin{matrix}\alpha_{2}q^{p},\beta_{2}\\ \gamma_{2}\end{matrix};q,y\right]z^{p},$ (14)
where $|x|<1$, $|y|<1$, $|z|<1$ and $_{r}\phi_{s}$ is the basic hypergeometric series defined by [19, p. 4, Eq. (1.2.22)]
$_{r}\phi_{s}\left[\begin{matrix}a_{1},\cdots,a_{r}\\ b_{1},\cdots,b_{s}\end{matrix};q,z\right]:=\sum_{n=0}^{\infty}\frac{(a_{1},\cdots,a_{r};q)_{n}}{(b_{1},\cdots,b_{s},q;q)_{n}}\Big((-1)^{n}q^{\binom{n}{2}}\Big)^{1+s-r}z^{n}.$
For simplicity, we write
$\widetilde{\Phi}_{K}[\alpha_{1},\alpha_{2},\alpha_{2},\beta_{1},\beta_{2},\beta_{1};\gamma_{1},\gamma_{2},\gamma_{3};q,x,y,z]$
$\hskip 14.22636pt=\Phi_{K}\big[q^{\alpha_{1}},q^{\alpha_{2}},q^{\alpha_{2}},q^{\beta_{1}},q^{\beta_{2}},q^{\beta_{1}};q^{\gamma_{1}},q^{\gamma_{2}},q^{\gamma_{3}};q,x,y,z\big]$ (15)
and
$_{r}\widetilde{\phi}_{s}\left[\begin{matrix}a_{1},\cdots,a_{r}\\ b_{1},\cdots,b_{s}\end{matrix};q,z\right]={}_{r}\phi_{s}\left[\begin{matrix}q^{a_{1}},\cdots,q^{a_{r}}\\ q^{b_{1}},\cdots,q^{b_{s}}\end{matrix};q,z\right].$
The only $q$-integral representation (of Bateman type) for $\Phi_{K}$ is given by Ernst [15, p. 10, Theorem 5]:
$\widetilde{\Phi}_{K}[\alpha_{1},\alpha_{2},\alpha_{2},\beta_{1},\beta_{2},\beta_{1};\gamma_{1},\gamma_{2},\gamma_{3};q,x,y,z]$
$\hskip 14.22636pt=\int_{[0,1]^{3}}\widetilde{\Phi}_{K}[\alpha_{1},\alpha_{2},\alpha_{2},\beta_{1},\beta_{2},\beta_{1};\nu_{1},\nu_{2},\nu_{3};q,xu,yv,zw]$
$\hskip 42.67912pt\cdot\mathrm{d}\mu_{\nu_{1},\gamma_{1}-\nu_{1}}(u;q)\mathrm{d}\mu_{\nu_{2},\gamma_{2}-\nu_{2}}(v;q)\mathrm{d}\mu_{\nu_{3},\gamma_{3}-\nu_{3}}(w;q).$ (16)
Here and in what follows, we define the $q$-analogue of the Dirichlet measure (3) as follows:
$\mathrm{d}\mu_{\alpha,\beta}(t;q):=\mathfrak{m}_{\alpha,\beta}(t;q)\mathrm{d}_{q}t,$ (17)
where
$\mathfrak{m}_{\alpha,\beta}(t;q):=\frac{\Gamma_{q}(\alpha+\beta)}{\Gamma_{q}(\alpha)\Gamma_{q}(\beta)}t^{\alpha-1}\frac{(tq;q)_{\infty}}{(tq^{\beta};q)_{\infty}},\penalty 10000\ \penalty 10000\ \penalty 10000\ \min\{\Re(\alpha),\Re(\beta)\}>0.$
Obviously, we have $\int_{0}^{1}\mathrm{d}\mu_{\alpha,\beta}(t;q)=1$ and $\lim_{q\rightarrow 1^{-}}\mathfrak{m}_{\alpha,\beta}(t;q)=\mathfrak{m}_{\alpha,\beta}(t)$.
Actually, in Section 4, we shall first prove a Joshi-Vyas type theorem and derive some useful corollaries from it. Then a discrete analogue for one of the corollaries is obtained. Finally, we present a $q$-analogue of Theorem 1.1.
In the next section, we will further introduce some results related to $q$-series.
2 Preliminaries
Gasper [18] discussed the $q$-analogues of Erdélyi’s integrals (5)–(7) and provided their discrete analogues. The integral (5) has the following $q$-analogue [18, p. 4, Eq. (1.13)]
${}_{2}\widetilde{\phi}_{1}\left[\begin{matrix}\alpha,\beta\\ \gamma\end{matrix};q,x\right]$ $=\int_{0}^{1}\frac{(xtq^{\alpha^{\prime}};q)_{\infty}}{(xt;q)_{\infty}}{}_{2}\widetilde{\phi}_{1}\left[\begin{matrix}\alpha-\alpha^{\prime},\beta\\ \lambda\end{matrix};q,xtq^{\alpha^{\prime}}\right]$
$\hskip 14.22636pt\cdot{}_{3}\phi_{2}\left[\begin{matrix}q^{\alpha^{\prime}},q^{\beta-\lambda},t^{-1}\\ q^{\gamma-\lambda},q/(xt)\end{matrix};q,q\right]\mathrm{d}\mu_{\lambda,\gamma-\lambda}(t;q),$ (18)
where $\Re(\gamma)>\Re(\lambda)>0$ and $\mathrm{d}\mu_{\alpha,\beta}(t;q)$ is defined in (17). Before stating the $q$-analogue of (7), we have to first introduce the $q$-analogue of the hypergeometric measure (8):
$\mathrm{d}\mu_{\alpha,\beta,\gamma,\eta}(t;q):=\mathfrak{m}_{\alpha,\beta,\gamma,\eta}(t;q)\mathrm{d}_{q}t,$ (19)
where $\min\{\Re(\eta),\Re(\gamma),\Re(\eta+\gamma-\alpha-\beta)\}>0$ and
$\mathfrak{m}_{\alpha,\beta,\gamma,\eta}(t;q)$ $=\frac{\Gamma_{q}(\eta+\gamma-\alpha)\Gamma_{q}(\eta+\gamma-\beta)}{\Gamma_{q}(\eta)\Gamma_{q}(\gamma)\Gamma_{q}(\eta+\gamma-\alpha-\beta)}$
$\hskip 28.45274pt\cdot t^{\eta-1}\frac{(tq;q)_{\infty}}{(tq^{\gamma};q)_{\infty}}{}_{3}\phi_{1}\left[\begin{matrix}q^{\alpha},q^{\beta},t^{-1}\\ q^{\gamma}\end{matrix};q,tq^{\gamma-\alpha-\beta}\right].$ (20)
We have $\int_{0}^{1}\mathrm{d}\mu_{\eta-\lambda,\gamma-\lambda,\gamma-\lambda+\eta-\nu,\nu}(t;q)=1$ and $\lim_{q\rightarrow 1^{-}}\mathfrak{m}_{\alpha,\beta,\gamma,\eta}(t;q)=\mathfrak{m}_{\alpha,\beta,\gamma,\eta}(t)$. Then the $q$-analogue of (7) is given by [18, p. 4, Eq. (1.14)]
${}_{2}\widetilde{\phi}_{1}\left[\begin{matrix}\alpha,\beta\\ \gamma\end{matrix};q,x\right]$ $=\int_{0}^{1}{}_{3}\widetilde{\phi}_{2}\left[\begin{matrix}\alpha,\beta,\eta\\ \lambda,\nu\end{matrix};q,xt\right]\mathrm{d}\mu_{\eta-\lambda,\gamma-\lambda,\gamma-\lambda+\eta-\nu,\nu}(t;q).$ (21)
where $\min\{\Re(\lambda),\Re(\nu),\Re(\gamma+\eta-\lambda-\nu)\}>0$. Joshi and Vyas [22] also derived some nice $q$-Erdélyi-type integrals, one of which is discussed in Section 4.
In the same paper [18], Gasper also found the following discrete $q$-analogue of (21):
${}_{3}\phi_{2}\left[\begin{matrix}\alpha,\beta,q^{-n}\\ \gamma,\delta\end{matrix};q,q\right]$ $=\frac{(q,\lambda;q)_{n}}{(\gamma,\mu;q)_{n}}\sum_{k=0}^{n}\frac{(\nu;q)_{k}(\gamma\mu/(\lambda\nu);q)_{n-k}}{(q;q)_{k}(q;q)_{n-k}}\nu^{n-k}$
$\cdot{}_{3}\phi_{2}\left[\begin{matrix}\mu/\lambda,\gamma/\lambda,q^{k-n}\\ \gamma\mu/(\lambda\nu),q^{1-n}/\lambda\end{matrix};q,q^{1-k}/\nu\right]{}_{4}\phi_{3}\left[\begin{matrix}\alpha,\beta,\mu,q^{-k}\\ \lambda,\nu,\delta\end{matrix};q,q\right].$ (22)
Finally, we also need the following $q$-hypergeometric function of three variables (see [9, p. 96] and [40, p. 35]):
$\phi^{(3)}[x,y,z]$ $\equiv\phi^{(3)}\left[\begin{matrix}(a)::(b);(b^{\prime});(b^{\prime\prime}):(c);(c^{\prime});(c^{\prime\prime})\\ (e)::(g);(g^{\prime});(g^{\prime\prime}):(h);(h^{\prime});(h^{\prime\prime})\end{matrix};q,x,y,z\right]$
$:=\sum_{m,n,p=0}^{\infty}\frac{\displaystyle\prod_{r=1}^{A}(a_{r};q)_{m+n+p}}{\displaystyle\prod_{r=1}^{E}(e_{r};q)_{m+n+p}}\cdot\frac{\displaystyle\prod_{r=1}^{B}(b_{r};q)_{m+n}\prod_{r=1}^{B^{\prime}}(b_{r}^{\prime};q)_{n+p}\prod_{r=1}^{B^{\prime\prime}}(b_{r}^{\prime\prime};q)_{p+m}}{\displaystyle\prod_{r=1}^{G}(g_{r};q)_{m+n}\prod_{r=1}^{G^{\prime}}(g_{r}^{\prime};q)_{n+p}\prod_{r=1}^{G^{\prime\prime}}(g_{r}^{\prime\prime};q)_{p+m}}$
$\hskip 14.22636pt\cdot\frac{\displaystyle\prod_{r=1}^{C}(c_{r};q)_{m}\prod_{r=1}^{C^{\prime}}(c_{r}^{\prime};q)_{n}\prod_{r=1}^{C^{\prime\prime}}(c_{r}^{\prime\prime};q)_{p}}{\displaystyle\prod_{r=1}^{H}(h_{r};q)_{m}\prod_{r=1}^{H^{\prime}}(h_{r}^{\prime};q)_{n}\prod_{r=1}^{H^{\prime\prime}}(h_{r}^{\prime\prime};q)_{p}}\cdot\frac{x^{m}}{(q;q)_{m}}\frac{y^{n}}{(q;q)_{n}}\frac{z^{p}}{(q;q)_{p}},$ (23)
where $|x|<1$, $|y|<1$ and $|z|<1$.
3 Theorem 1.1 revisited
3.1 An alternative proof of Theorem 1.1
For convenience, let $I$ denote the triple integral in (1.1). By (10), the first $F_{K}$-function in the integrand can be expanded as
$F_{K}[\alpha_{1},\alpha_{2}-\eta_{2},\alpha_{2}-\eta_{2},\beta_{1}-\lambda_{3},\beta_{2},\beta_{1}-\lambda_{3};\alpha_{1}-\lambda_{1}+\eta_{1},\beta_{2}-\lambda_{2}+\mu_{2},\beta_{1}-\lambda_{3};ux,vy,wz]$
$\hskip 28.45274pt=\sum_{m=0}^{\infty}\frac{(\alpha_{2}-\eta_{2})_{m}}{m!}{}_{2}F_{1}\left[\begin{matrix}\beta_{1}-\lambda_{3}+m,\alpha_{1}\\ \alpha_{1}-\lambda_{1}+\eta_{1}\end{matrix};ux\right]{}_{2}F_{1}\left[\begin{matrix}\alpha_{2}-\eta_{2}+m,\beta_{2}\\ \beta_{2}-\lambda_{2}+\mu_{2}\end{matrix};vy\right](wz)^{m}.$ (24)
Since $\mathbb{D}_{K}$ is a Reinhardt domain, we have immediately $(ux,vy,wz)\in\mathbb{D}_{K}$ $(u,v,w\in(0,1))$ and thus the region of convergence of the series in (3.1) is clear. For the second $F_{K}$-function in the integrand, we have
$F_{K}\left[\lambda_{1}-\eta_{1},\eta_{2},\eta_{2},\lambda_{3},\lambda_{2}-\mu_{2},\lambda_{3};\lambda_{1},\lambda_{2},\lambda_{3};\frac{(1-u)x}{1-ux},\frac{(1-v)y}{1-vy},\frac{wz}{(1-ux)(1-vy)}\right]$
$\hskip 14.22636pt=\sum_{n=0}^{\infty}\frac{(\eta_{2})_{n}}{n!}{}_{2}F_{1}\left[\begin{matrix}\lambda_{3}+n,\lambda_{1}-\eta_{1}\\ \lambda_{1}\end{matrix};\frac{(1-u)x}{1-ux}\right]$
$\hskip 42.67912pt\cdot{}_{2}F_{1}\left[\begin{matrix}\eta_{2}+n,\lambda_{2}-\mu_{2}\\ \lambda_{2}\end{matrix};\frac{(1-v)y}{1-vy}\right]\cdot\left[\frac{wz}{(1-ux)(1-vy)}\right]^{n}.$ (25)
Note that
$\left|\frac{(1-u)x}{1-ux}\right|<\frac{1-u}{1-u|x|}<1,\penalty 10000\ \penalty 10000\ \penalty 10000\ \left|\frac{(1-v)x}{1-vx}\right|<\frac{1-v}{1-v|x|}<1$
and
$\frac{|\frac{wz}{(1-ux)(1-vy)}|}{\big(1-\big|\frac{(1-u)x}{1-ux}\big|\big)\big(1-\big|\frac{(1-v)y}{1-vy}\big|\big)}$ $=\frac{w|z|}{(|1-ux|-(1-u)|x|)(|1-vy|-(1-v)|y|)}$
$\leq\frac{|z|}{(1-|x|)(1-|y|)}<1,$
namely,
$\left(\frac{(1-u)x}{1-ux},\frac{(1-v)y}{1-vy},\frac{wz}{(1-ux)(1-vy)}\right)\in\mathbb{D}_{K}.$
Thus the region of convergence of the series in (3.1) is also fully characterized.
Now we have
$I$ $=\sum_{m=0}^{\infty}\frac{(\alpha_{2}-\eta_{2})_{m}}{m!}z^{m}\sum_{n=0}^{\infty}\frac{(\eta_{2})_{n}}{n!}z^{n}\int_{0}^{1}w^{m+n}\mathrm{d}\mu_{\beta_{1},\gamma_{3}-\beta_{1}}(w)$
$\hskip 14.22636pt\cdot\int_{0}^{1}(1-ux)^{-\lambda_{3}-n}{}_{2}F_{1}\left[\begin{matrix}\beta_{1}-\lambda_{3}+m,\alpha_{1}\\ \alpha_{1}-\lambda_{1}+\eta_{1}\end{matrix};ux\right]{}_{2}F_{1}\left[\begin{matrix}\lambda_{3}+n,\lambda_{1}-\eta_{1}\\ \lambda_{1}\end{matrix};\frac{(1-u)x}{1-ux}\right]\mathrm{d}\mu_{\alpha_{1}-\lambda_{1}+\eta_{1},\lambda_{1}}(u)$
$\hskip 14.22636pt\cdot\int_{0}^{1}(1-vy)^{-\eta_{2}-n}{}_{2}F_{1}\left[\begin{matrix}\alpha_{2}-\eta_{2}+m,\beta_{2}\\ \beta_{2}-\lambda_{2}+\mu_{2}\end{matrix};vy\right]{}_{2}F_{1}\left[\begin{matrix}\eta_{2}+n,\lambda_{2}-\mu_{2}\\ \lambda_{2}\end{matrix};\frac{(1-v)y}{1-vy}\right]\mathrm{d}\mu_{\beta_{2}-\lambda_{2}+\mu_{2},\lambda_{2}}(v).$ (26)
Note that
$\int_{0}^{1}w^{m+n}\mathrm{d}\mu_{\beta_{1},\gamma_{3}-\beta_{1}}(w)=\frac{(\beta_{1})_{m+n}}{(\gamma_{3})_{m+n}},$ (27)
where $\Re(\gamma_{3})>\Re(\beta_{1})>0$. By letting $\alpha^{\prime}\rightarrow\lambda_{3}+n$, $\beta\rightarrow\alpha_{1}$, $\lambda\rightarrow\alpha_{1}-\lambda_{1}+\eta_{1}$, $\alpha\rightarrow\beta_{1}+m+n$ and $\gamma\rightarrow\alpha_{1}+\eta_{1}$ in the Erdélyi integral (5), we obtain
$\int_{0}^{1}(1-ux)^{-\lambda_{3}-n}{}_{2}F_{1}\left[\begin{matrix}\beta_{1}-\lambda_{3}+m,\alpha_{1}\\ \alpha_{1}-\lambda_{1}+\eta_{1}\end{matrix};ux\right]$
$\hskip 14.22636pt\cdot{}_{2}F_{1}\left[\begin{matrix}\lambda_{3}+n,\lambda_{1}-\eta_{1}\\ \lambda_{1}\end{matrix};\frac{(1-u)x}{1-ux}\right]\mathrm{d}\mu_{\alpha_{1}-\lambda_{1}+\eta_{1},\lambda_{1}}(u)={}_{2}F_{1}\left[\begin{matrix}\beta_{1}+m+n,\alpha_{1}\\ \alpha_{1}+\eta_{1}\end{matrix};x\right],$ (28)
where $\Re(\alpha_{1}+\eta_{1})>\Re(\lambda_{1})>0$. By letting $\alpha^{\prime}\rightarrow\eta_{2}+n$, $\beta\rightarrow\beta_{2}$, $\lambda\rightarrow\beta_{2}-\lambda_{2}+\mu_{2}$, $\alpha\rightarrow\alpha_{2}+m+n$ and $\gamma\rightarrow\beta_{2}+\mu_{2}$ in the Erdélyi integral (5), we obtain
$\int_{0}^{1}(1-vy)^{-\eta_{2}-n}{}_{2}F_{1}\left[\begin{matrix}\alpha_{2}-\eta_{2}+m,\beta_{2}\\ \beta_{2}-\lambda_{2}+\mu_{2}\end{matrix};vy\right]$
$\hskip 14.22636pt\cdot{}_{2}F_{1}\left[\begin{matrix}\eta_{2}+n,\lambda_{2}-\mu_{2}\\ \lambda_{2}\end{matrix};\frac{(1-v)y}{1-vy}\right]\mathrm{d}\mu_{\beta_{2}-\lambda_{2}+\mu_{2},\lambda_{2}}(v)={}_{2}F_{1}\left[\begin{matrix}\alpha_{2}+m+n,\beta_{2}\\ \beta_{2}+\mu_{2}\end{matrix};y\right],$ (29)
where $\Re(\beta_{2}+\mu_{2})>\Re(\lambda_{2})>0$. Substituting (27), (3.1) and (3.1) in (3.1) gives
$I$ $=\sum_{m,n=0}^{\infty}\frac{(\beta_{1})_{m+n}(\alpha_{2}-\eta_{2})_{m}(\eta_{2})_{n}}{(\gamma_{3})_{m+n}m!n!}{}_{2}F_{1}\left[\begin{matrix}\beta_{1}+m+n,\alpha_{1}\\ \alpha_{1}+\eta_{1}\end{matrix};x\right]{}_{2}F_{1}\left[\begin{matrix}\alpha_{2}+m+n,\beta_{2}\\ \beta_{2}+\mu_{2}\end{matrix};y\right]z^{m+n}$
$=\sum_{m=0}^{\infty}\frac{(\beta_{1})_{m}}{(\gamma_{3})_{m}}{}_{2}F_{1}\left[\begin{matrix}\beta_{1}+m,\alpha_{1}\\ \alpha_{1}+\eta_{1}\end{matrix};x\right]{}_{2}F_{1}\left[\begin{matrix}\alpha_{2}+m,\beta_{2}\\ \beta_{2}+\mu_{2}\end{matrix};y\right]z^{m}\sum_{n=0}^{m}\frac{(\alpha_{2}-\eta_{2})_{m-n}(\eta_{2})_{n}}{(m-n)!n!}.$
The result (1.1) now follows by applying the Chu-Vandermonde identity
$(a+b)_{m}=\sum_{n=0}^{m}\binom{m}{n}(a)_{m-n}(b)_{n}.$
Remark 3.1.
Joshi and Vyas [21] have shown that the Erdélyi integrals (5)-(7) can be derived by using the series manipulation technique. Thus, inspired by the proof presented in this section, we can also write down a proof using only the series manipulation technique. However, such a proof is quite complicated, so we choose to omit it here.
3.2 A curious integral representation for $F_{2}$
In this subsection, we derive a neat and interesting integral from Theorem 1.1. This integral is novel and has not been mentioned in the existing literature.
Theorem 3.2.
Let $\Re(c_{1})>\Re(d_{1})>0$ and $\Re(c_{2})>\Re(b_{2})>0$. Then we have
$F_{2}[a_{1},b_{1},b_{2};c_{1},c_{2};y,z]$ $=\int_{0}^{1}\int_{0}^{1}(1-wz-vy)^{-a_{1}}{}_{2}F_{1}\left[\begin{matrix}a_{1}-a_{2},c_{1}-b_{1}-d_{1}\\ c_{1}-d_{1}\end{matrix};\frac{vy}{vy+wz-1}\right]$
$\hskip 14.22636pt\cdot{}_{2}F_{1}\left[\begin{matrix}a_{2},b_{1}+d_{1}-c_{1}\\ d_{1}\end{matrix};\frac{(1-v)y}{1-vy-wz}\right]\mathrm{d}\mu_{c_{1}-d_{1},d_{1}}(v)\mathrm{d}\mu_{b_{2},c_{2}-b_{2}}(w).$ (30)
Proof.
Letting $x=0$ in (1.1), it reduces to
$F_{2}[\alpha_{2},\beta_{2},\beta_{1};\beta_{2}+\mu_{2},\gamma_{3};y,z]$
$\hskip 14.22636pt=\int_{0}^{1}\int_{0}^{1}(1-vy)^{-\eta_{2}}F_{2}[\alpha_{2}-\eta_{2},\beta_{2},\beta_{1}-\lambda_{3};\beta_{2}-\lambda_{2}+\mu_{2},\beta_{1}-\lambda_{3};vy,wz]$
$\hskip 42.67912pt\cdot F_{2}\left[\eta_{2},\lambda_{2}-\mu_{2},\lambda_{3};\lambda_{2},\lambda_{3};\frac{(1-v)y}{1-vy},\frac{wz}{1-vy}\right]\mathrm{d}\mu_{\beta_{2}-\lambda_{2}+\mu_{2},\lambda_{2}}(v)\mathrm{d}\mu_{\beta_{1},\gamma_{3}-\beta_{1}}(w).$ (31)
Next, letting $\alpha_{2}\rightarrow a_{1}$, $\beta_{2}\rightarrow b_{1}$, $\beta_{1}\rightarrow b_{2}$, $\mu_{2}\rightarrow c_{1}-b_{1}$, $\gamma_{3}\rightarrow c_{2}$, $\lambda_{2}\rightarrow d_{1}$, $\lambda_{3}\rightarrow d_{2}$ and $\eta_{2}\rightarrow a_{2}$ in (3.2) gives
$F_{2}[a_{1},b_{1},b_{2};c_{1},c_{2};y,z]$
$\hskip 14.22636pt=\int_{0}^{1}\int_{0}^{1}(1-vy)^{-a_{2}}F_{2}[a_{1}-a_{2},b_{1},b_{2}-d_{2};c_{1}-d_{1},b_{2}-d_{2};vy,wz]$
$\hskip 42.67912pt\cdot F_{2}\left[a_{2},d_{1}-c_{1}+b_{1},d_{2};d_{1},d_{2};\frac{(1-v)y}{1-vy},\frac{wz}{1-vy}\right]\mathrm{d}\mu_{c_{1}-d_{1},d_{1}}(v)\mathrm{d}\mu_{b_{2},c_{2}-b_{2}}(w).$ (32)
Then, by using [48, p. 306, Eq. (109)]
$F_{2}[a,b,b^{\prime};c,b^{\prime};y,z]=(1-z)^{-a}{}_{2}F_{1}\left[\begin{matrix}a,b\\ c\end{matrix};\frac{y}{1-z}\right].$ (33)
we have
$(1-vy)^{-a_{2}}F_{2}\left[a_{2},d_{1}-c_{1}+b_{1},d_{2};d_{1},d_{2};\frac{(1-v)y}{1-vy},\frac{wz}{1-vy}\right]$
$\hskip 28.45274pt=(1-vy-wz)^{-a_{2}}{}_{2}F_{1}\left[\begin{matrix}a_{2},b_{1}+d_{1}-c_{1}\\ d_{1}\end{matrix};\frac{(1-v)y}{1-vy-wz}\right].$ (34)
By using (33) and the Pfaff transformation [48, p. 300, Eq. (79)], we have
$F_{2}[a_{1}-a_{2},b_{1},b_{2}-d_{2};c_{1}-d_{1},b_{2}-d_{2};vy,wz]$
$\hskip 28.45274pt=(1-wz)^{-a_{1}+a_{2}}{}_{2}F_{1}\left[\begin{matrix}a_{1}-a_{2},b_{1}\\ c_{1}-d_{1}\end{matrix};\frac{vy}{1-wz}\right]$
$\hskip 28.45274pt=(1-wz-vy)^{-a_{1}+a_{2}}{}_{2}F_{1}\left[\begin{matrix}a_{1}-a_{2},c_{1}-b_{1}-d_{1}\\ c_{1}-d_{1}\end{matrix};\frac{vy}{vy+wz-1}\right].$ (35)
Substituting (3.2) and (3.2) into (3.2) leads us to (3.2). ∎
Remark 3.3.
We can prove (3.2) directly by using the method described in Section 3.1. In addition, the integral (3.2) can also be derived from Manocha’s integral [34, p. 239, Eq. (13)]:
$F_{2}[a,b,c;d,e;y,z]=\int_{0}^{1}\int_{0}^{1}(1-vy-wz)^{-a^{\prime}}F_{2}[a-a^{\prime},b,c;\lambda,\eta;yv,zw]$
$\hskip 28.45274pt\cdot F_{2}\left[a^{\prime},b-\lambda,c-\eta;d-\lambda,e-\eta;\frac{(1-v)y}{1-vx-wz},\frac{(1-w)z}{1-vy-wz}\right]\mathrm{d}\mu_{\lambda,d-\lambda}(v)\mathrm{d}\mu_{\eta,e-\eta}(w),$ (36)
where $\Re(d)>\Re(\lambda)>0$ and $\Re(e)>\Re(\eta)>0$. In fact, if we set $\eta\rightarrow c$ in (3.3) and then use the two transformation formulas as in the proof above, we obtain
$F_{2}[a,b,c;d,e;y,z]=\int_{0}^{1}\int_{0}^{1}(1-vy-wz)^{-a}{}_{2}F_{1}\left[\begin{matrix}a-a^{\prime},\lambda-b\\ \lambda\end{matrix};\frac{vy}{vy+wz-1}\right]$
$\hskip 28.45274pt\cdot{}_{2}F_{1}\left[\begin{matrix}a^{\prime},b-\lambda\\ d-\lambda\end{matrix};\frac{(1-v)y}{1-vy-wz}\right]\mathrm{d}\mu_{\lambda,d-\lambda}(v)\mathrm{d}\mu_{c,e-c}(w).$ (37)
Letting further $\lambda\rightarrow c_{1}-d_{1}$, $b\rightarrow b_{1}$, $\rightarrow a_{1}, aa$$^{\prime}\rightarrow a_{2}$, $c\rightarrow b_{2}$, $d\rightarrow c_{1}$ and $e\rightarrow c_{2}$ in (3.3) leads us again to the integral (3.2).
3.3 A further extension of Theorem 1.1
As usual, the discrete convolution product of two sequences $\mathbf{a}:=a(m,n)$ and $\mathbf{b}:=b(m,n)$ is defined by
$(\mathbf{a}\star\mathbf{b})(m,n):=\sum_{i=0}^{m}\sum_{j=0}^{n}a(m-i,n-j)b(i,j).$
Let us consider the function of the following form:
$\mathcal{F}^{\mathbf{a}}\left[\begin{matrix}\alpha_{1},\beta_{1}:\alpha_{2},\beta_{2}\\ \gamma_{1}:\gamma_{2}\end{matrix};x_{1},x_{2},x_{3},x_{4}\right]$ $:=\sum_{m,n=0}^{\infty}a(m,n){}_{2}F_{1}\left[\begin{matrix}\alpha_{1}+m,\beta_{1}\\ \gamma_{1}\end{matrix};x_{1}\right]$
$\hskip 14.22636pt\cdot{}_{2}F_{1}\left[\begin{matrix}\alpha_{2}+n,\beta_{2}\\ \gamma_{2}\end{matrix};x_{2}\right]x_{3}^{m}x_{4}^{n}.$ (38)
Functions of this type appear frequently in many situations and have properties similar to Saran’s $F_{K}$-function.
If we take
$a(m,n)=\begin{cases}\displaystyle\frac{(\alpha_{1})_{n}(\alpha_{2})_{n}}{n!(\gamma_{3})_{n}},&m=n,\\ 0,&m\neq n,\end{cases}$
then the function (3.3) reduces to Saran’s function:
$F_{K}\left[\beta_{1},\alpha_{2},\alpha_{2},\alpha_{1},\beta_{2},\alpha_{1};\gamma_{1},\gamma_{2},\gamma_{3};x_{1},x_{2},x_{3}x_{4}\right].$
Another significant example originates in physics. In 2018, Belitsky [4, p. 69] generalized Saran’s $F_{K}$-function of three variables to $L$ variables, i.e.,
$F_{K}\left[\begin{matrix}a_{1},b_{1},\cdots,b_{L-1},a_{2}\\ c_{1},\cdots,c_{L}\end{matrix};z_{1},\cdots,z_{L}\right]$
$\hskip 28.45274pt:=\sum_{n_{1},\cdots,n_{L}=0}^{\infty}\frac{\left(a_{1}\right)_{n_{1}}\left(b_{1}\right)_{n_{1}+n_{2}}\cdots\left(b_{L-1}\right)_{n_{L-1}+n_{L}}\left(a_{2}\right)_{n_{L}}}{\left(c_{1}\right)_{n_{1}}\cdots\left(c_{L}\right)_{n_{L}}}\frac{z_{1}^{n_{1}}}{n_{1}!}\cdots\frac{z_{L}^{n_{L}}}{n_{L}!}.$ (39)
Then Rosenhaus [42] employed the function (3.3) in his study of $n$-point conformal blocks in the comb channel. More recently, Ferrando et al. [16] used a rescaled version of this function in evaluating the conformal integral associated with the $n$-point conformal partial wave in the comb channel. Taking $L=4$ in (3.3), we obtain
$F_{K}\left[\begin{matrix}a_{1},b_{1},b_{2},b_{3},a_{2}\\ c_{1},c_{2},c_{3},c_{4}\end{matrix};z_{1},z_{2},z_{3},z_{4}\right]$
$\hskip 14.22636pt=\sum_{n_{1},n_{2},n_{3},n_{4}=0}^{\infty}\frac{(a_{1})_{n_{1}}(b_{1})_{n_{1}+n_{2}}(b_{2})_{n_{2}+n_{3}}(b_{3})_{n_{3}+n_{4}}(a_{2})_{n_{4}}}{(c_{1})_{n_{1}}(c_{2})_{n_{2}}(c_{3})_{n_{3}}(c_{4})_{n_{4}}}\frac{z_{1}^{n_{1}}}{n_{1}!}\frac{z_{2}^{n_{2}}}{n_{2}!}\frac{z_{3}^{n_{3}}}{n_{3}!}\frac{z_{4}^{n_{4}}}{n_{4}!}$
$\hskip 14.22636pt=\sum_{n_{2},n_{3}=0}^{\infty}\frac{(b_{1})_{n_{2}}(b_{2})_{n_{2}+n_{3}}(b_{3})_{n_{3}}}{(c_{2})_{n_{2}}(c_{3})_{n_{3}}n_{2}!n_{3}!}{}_{2}F_{1}\left[\begin{matrix}a_{1},b_{1}+n_{2}\\ c_{1}\end{matrix};z_{1}\right]{}_{2}F_{1}\left[\begin{matrix}a_{2},b_{3}+n_{3}\\ c_{4}\end{matrix};z_{4}\right]z_{2}^{n_{2}}z_{3}^{n_{3}},$ (40)
which also has the form of (3.3). The region of convergence of (3.3) is determined by
$|z_{1}|<1,\penalty 10000\ |z_{4}|<1\penalty 10000\ \penalty 10000\ \penalty 10000\ \text{and}\penalty 10000\ \penalty 10000\ \penalty 10000\ \frac{|z_{2}|}{1-|z_{1}|}+\frac{|z_{3}|}{1-|z_{4}|}<1.$
By using the method from Section 3.1, we have the following theorem.
Theorem 3.4.
Let $\Re(\tau_{j})>\Re(\gamma_{j})>0$ $(j=1,\cdots,4)$. Then we have
$\mathcal{F}^{\mathbf{c}}\left[\begin{matrix}\alpha_{1}+\lambda_{1},\beta_{1}:\alpha_{2}+\lambda_{2},\beta_{2}\\ \tau_{1}:\tau_{2}\end{matrix};x_{1},x_{2},x_{3},x_{4}\right]=\int_{0}^{1}\int_{0}^{1}\int_{0}^{1}\int_{0}^{1}(1-u_{1}x_{1})^{-\lambda_{1}}(1-u_{2}x_{2})^{-\lambda_{2}}$
$\hskip 14.22636pt\cdot\mathcal{F}^{\mathbf{a}}\left[\begin{matrix}\alpha_{1},\beta_{1}:\alpha_{2},\beta_{2}\\ \gamma_{1}:\gamma_{2}\end{matrix};x_{1}u_{1},x_{2}u_{2},x_{3}u_{3},x_{4}u_{4}\right]$
$\hskip 14.22636pt\cdot\mathcal{F}^{\mathbf{b}}\left[\begin{matrix}\lambda_{1},\beta_{1}-\gamma_{1}:\lambda_{2},\beta_{2}-\gamma_{2}\\ \tau_{1}-\gamma_{1}:\tau_{2}-\gamma_{2}\end{matrix};\frac{(1-u_{1})x_{1}}{1-u_{1}x_{1}},\frac{(1-u_{2})x_{2}}{1-u_{2}x_{2}},\frac{u_{3}x_{3}}{1-u_{1}x_{1}},\frac{u_{4}x_{4}}{1-u_{2}x_{2}}\right]\prod_{j=1}^{4}\mathrm{d}\mu_{\gamma_{j},\tau_{j}-\gamma_{j}}(u_{j}),$ (41)
where
$\mathbf{c}(m,n)=\frac{(\gamma_{3})_{m}}{(\tau_{3})_{m}}\frac{(\gamma_{4})_{n}}{(\tau_{4})_{n}}(\mathbf{a}\star\mathbf{b})(m,n).$
Proof.
Since the proof is very similar to the one given in Section 3.1, we omit the details and retain only key steps. Let us denote the right-hand side of (3.4) by $I$. Then, by using (3.3), we have
$I$ $=\sum_{m_{1},m_{2}=0}^{\infty}\sum_{n_{1},n_{2}=0}^{\infty}a(m_{1},m_{2})b(n_{1},n_{2})x_{3}^{m_{1}+n_{1}}x_{4}^{m_{2}+n_{2}}$
$\hskip 14.22636pt\cdot\int_{0}^{1}(1-u_{1}x_{1})^{-\lambda_{1}-n_{1}}{}_{2}F_{1}\left[\begin{matrix}\alpha_{1}+m_{1},\beta_{1}\\ \gamma_{1}\end{matrix};x_{1}u_{1}\right]{}_{2}F_{1}\left[\begin{matrix}\lambda_{1}+n_{1},\beta_{1}-\gamma_{1}\\ \tau_{1}-\gamma_{1}\end{matrix};\frac{(1-u_{1})x_{1}}{1-u_{1}x_{1}}\right]\mathrm{d}\mu_{\gamma_{1},\tau_{1}-\gamma_{1}}(u_{1})$
$\hskip 14.22636pt\cdot\int_{0}^{1}(1-u_{2}x_{2})^{-\lambda_{2}-n_{2}}{}_{2}F_{1}\left[\begin{matrix}\alpha_{2}+m_{2},\beta_{2}\\ \gamma_{2}\end{matrix};x_{2}u_{2}\right]{}_{2}F_{1}\left[\begin{matrix}\lambda_{2}+n_{2},\beta_{2}-\gamma_{2}\\ \tau_{2}-\gamma_{2}\end{matrix};\frac{(1-u_{2})x_{2}}{1-u_{2}x_{2}}\right]\mathrm{d}\mu_{\gamma_{2},\tau_{2}-\gamma_{2}}(u_{2})$
$\hskip 14.22636pt\cdot\int_{0}^{1}u_{3}^{m_{1}+n_{1}}\mathrm{d}\mu_{\gamma_{3},\tau_{3}-\gamma_{3}}(u_{3})\int_{0}^{1}u_{4}^{m_{2}+n_{2}}\mathrm{d}\mu_{\gamma_{4},\tau_{4}-\gamma_{4}}(u_{4}).$ (42)
By using (5) and (27), we obtain
$I$ $=\sum_{m_{1},m_{2}=0}^{\infty}\sum_{n_{1},n_{2}=0}^{\infty}a(m_{1},m_{2})b(n_{1},n_{2})x_{3}^{m_{1}+n_{1}}x_{4}^{m_{2}+n_{2}}\frac{(\gamma_{3})_{m_{1}+n_{1}}}{(\tau_{3})_{m_{1}+n_{1}}}\frac{(\gamma_{4})_{m_{2}+n_{2}}}{(\tau_{4})_{m_{2}+n_{2}}}$
$\hskip 28.45274pt\cdot{}_{2}F_{1}\left[\begin{matrix}\alpha_{1}+\lambda_{1}+m_{1}+n_{1},\beta_{1}\\ \tau_{1}\end{matrix};x_{1}\right]{}_{2}F_{1}\left[\begin{matrix}\alpha_{2}+\lambda_{2}+m_{2}+n_{2},\beta_{2}\\ \tau_{2}\end{matrix};x_{2}\right],$
which, through the same series manipulation as in Section 3.1, leads us to the desired result. ∎
4 Some $q$-Erdélyi-type integrals
4.1 A Joshi-Vyas type theorem
Inspired by the formula [21, p. 133] of Joshi and Vyas and its useful $q$-analogue [22, p. 646], we establish–using the measure in (19)–the following theorem of Joshi-Vyas type.
Theorem 4.1.
Let $\Phi(z_{1},\cdots,z_{k})$ be defined by
$\Phi(z_{1},\cdots,z_{k}):=\sum_{n_{1},\cdots,n_{k}=0}^{\infty}c(n_{1},\cdots,n_{k})z_{1}^{n_{1}}\cdots z_{k}^{n_{k}}$
where $\{c(n_{1},\cdots,n_{k})\}$ is a suitable sequence so that the series converges absolutely and uniformly in a certain region. Then
$\sum_{n_{1},\cdots,n_{k}=0}^{\infty}c(n_{1},\cdots,n_{k})\prod_{j=1}^{k}\frac{(q^{\nu_{j}},q^{\lambda_{j}};q)_{n_{j}}}{(q^{\gamma_{j}},q^{\eta_{j}};q)_{n_{j}}}z_{j}^{n_{j}}$
$\hskip 14.22636pt=\int_{[0,1]^{k}}\Phi(z_{1}t_{1},\cdots,z_{k}t_{k})\prod_{j=1}^{k}\mathrm{d}\mu_{\eta_{j}-\lambda_{j},\gamma_{j}-\lambda_{j},\gamma_{j}-\lambda_{j}+\eta_{j}-\nu_{j},\nu_{j}}(t_{j};q).$ (43)
Proof.
The result follows by the moment formula
$\int_{0}^{1}t^{\ell}\mathrm{d}\mu_{\eta-\lambda,\gamma-\lambda,\gamma-\lambda+\eta-\nu,\nu}(t;q)=\frac{(q^{\nu},q^{\lambda};q)_{\ell}}{(q^{\gamma},q^{\eta};q)_{\ell}},$
which can be proved by evaluating the $q$-integral directly. This process, although a little complicated, is quite straightforward; hence we decide to omit it here. ∎
Remark 4.2.
When $k=1$, the formula (4.1) reduces to the result obtained by Joshi and Vyas. It can also be considered as Ernst’s $q$-analogue for Exton’s result (see [14, p. 189, Theorem 3.1]).
Letting $k=3$ and
$c(m,n,p)=(q^{\alpha_{2}};q)_{n+p}(q^{\beta_{1}};q)_{p+m}\frac{(q^{\alpha_{1}},q^{\eta_{1}};q)_{m}(q^{\beta_{2}},q^{\eta_{2}};q)_{n}(q^{\eta_{3}};q)_{p}}{(q^{\nu_{1}},q^{\lambda_{1}},q;q)_{m}(q^{\nu_{2}},q^{\lambda_{2}},q;q)_{n}(q^{\lambda_{3}},q^{\eta_{3}},q;q)_{p}}$
in Theorem 4.1 gives the following corollary.
Corollary 4.3.
When $\min\{\Re(\gamma_{i}+\eta_{i}-\lambda_{i}-\nu_{i}),\Re(\lambda_{i}),\Re(\nu_{i})\}>0$ $(i=1,2,3)$, we have
$\widetilde{\Phi}_{K}[\alpha_{1},\alpha_{2},\alpha_{2},\beta_{1},\beta_{2},\beta_{1},\gamma_{1},\gamma_{2},\gamma_{3};q,x,y,z]$
$\hskip 14.22636pt=\int_{[0,1]^{3}}\widetilde{\phi}^{(3)}\left[\begin{matrix}-::-;\alpha_{2};\beta_{1}:\alpha_{1},\eta_{1};\beta_{2},\eta_{2};\eta_{3}\\ -::-;-;-:\nu_{1},\lambda_{1};\nu_{2},\lambda_{2};\nu_{3},\lambda_{3}\end{matrix};q,xt_{1},yt_{2},zt_{3}\right]$
$\hskip 42.67912pt\cdot\prod_{j=1}^{3}\mathrm{d}\mu_{\eta_{j}-\lambda_{j},\gamma_{j}-\lambda_{j},\gamma_{j}-\lambda_{j}+\eta_{j}-\nu_{j},\nu_{j}}(t_{j};q),$ (44)
where $\phi^{(3)}$ is defined by (2).
Remark 4.4.
This integral (4.3) is clearly a $q$-analogue of the integral by Luo et al. [32, Theorem 1]. In particular, when $x=0$, (4.3) reduces to
$\widetilde{\Phi}_{2}[\alpha_{2},\beta_{2},\beta_{1};\gamma_{2},\gamma_{3};q,y,z]$ $=\int_{[0,1]^{2}}\widetilde{\Phi}_{0:2;2}^{1:2;2}\left[\begin{matrix}\alpha_{2}:\beta_{2},\eta_{2};\beta_{1},\eta_{3}\\ -:\nu_{2},\lambda_{2};\nu_{3},\lambda_{3}\end{matrix};q,yt_{2},zt_{3}\right]$
$\hskip 42.67912pt\cdot\prod_{j=2}^{3}\mathrm{d}\mu_{\eta_{j}-\lambda_{j},\gamma_{j}-\lambda_{j},\gamma_{j}-\lambda_{j}+\eta_{j}-\nu_{j},\nu_{j}}(t_{j};q),$ (45)
which is the $q$-analogue of Koschmieder’s integral representation for $F_{2}$ ([32, Corollary 1]). Here, $\Phi_{2}$ and $\Phi_{0:2;2}^{1:2;2}$ denote the $q$-Appell function [19, p. 283, Eq. (10.2.6)] and the $q$-Kampé de Fériet function [48, p. 349], respectively.
Letting $\lambda_{1}=\alpha_{1}$, $\lambda_{2}=\beta_{2}$ and $\lambda_{3}=\eta_{3}$ in (4.3), we obtain the following corollary.
Corollary 4.5.
When $\min\{\Re(\gamma_{1}+\eta_{1}-\alpha_{1}-\nu_{1}),\Re(\alpha_{1}),\Re(\nu_{1})\}>0$, $\min\{\Re(\gamma_{2}+\eta_{2}-\beta_{2}-\nu_{2}),\Re(\beta_{2}),\Re(\nu_{2})\}>0$ and $\Re(\gamma_{3})>\Re(\nu_{3})>0$, we have
$\widetilde{\Phi}_{K}[\alpha_{1},\alpha_{2},\alpha_{2},\beta_{1},\beta_{2},\beta_{1},\gamma_{1},\gamma_{2},\gamma_{3};q,x,y,z]$
$\hskip 14.22636pt=\int_{[0,1]^{3}}\widetilde{\Phi}_{K}[\eta_{1},\alpha_{2},\alpha_{2},\beta_{1},\eta_{2},\beta_{1};\nu_{1},\nu_{2},\nu_{3};q,xt_{1},yt_{2},zt_{3}]$
$\hskip 34.14322pt\cdot\mathrm{d}\mu_{\eta_{1}-\alpha_{1},\gamma_{1}-\alpha_{1},\gamma_{1}-\alpha_{1}+\eta_{1}-\nu_{1},\nu_{1}}(t_{1};q)\mathrm{d}\mu_{\eta_{2}-\beta_{2},\gamma_{2}-\beta_{2},\gamma_{2}-\beta_{2}+\eta_{2}-\nu_{2},\nu_{2}}(t_{2};q)\mathrm{d}\mu_{\nu_{3},\gamma_{3}-\nu_{3}}(t_{3};q).$ (46)
Remark 4.6.
This integral is the $q$-analogue of the integral by Luo and Raina [31, p. 14, Theorem 4.1].
In the next subsection, we shall show that Corollary 4.5 has a nice discrete analogue.
4.2 A discrete $q$-analogue of Corollary 4.5
Based on Gasper’s formula (2), we obtain the following results.
Theorem 4.7.
$\phi^{(3)}\left[\begin{matrix}-::-;q^{\alpha_{2}};q^{\beta_{1}}:q^{\alpha_{1}},q^{-r};q^{\beta_{2}},q^{-s};q^{-t}\\ -::-;-;-:q^{\gamma_{1}},\delta_{1};q^{\gamma_{2}},\delta_{2};q^{\gamma_{3}},\delta_{3}\end{matrix};q,q,q,q\right]$
$=\sum_{i=0}^{r}\sum_{j=0}^{s}\sum_{k=0}^{t}w_{1}(i,r;q)w_{2}(j,s;q)w_{3}(k,t;q)$
$\hskip 14.22636pt\cdot\phi^{(3)}\left[\begin{matrix}-::-;q^{\alpha_{2}};q^{\beta_{1}}:q^{\lambda_{1}},q^{-i};q^{\lambda_{2}},q^{-j};q^{-k}\\ -::-;-;-:q^{\mu_{1}},\delta_{1};q^{\mu_{2}},\delta_{2};q^{\mu_{3}},\delta_{3}\end{matrix};q,q,q,q\right],$ (47)
where
$w_{1}(i,r;q)$ $:=\frac{(q^{\alpha_{1}},q;q)_{r}}{(q^{\gamma_{1}},q^{\lambda_{1}};q)_{r}}\frac{(q^{\gamma_{1}+\lambda_{1}-\alpha_{1}-\mu_{1}};q)_{r-i}}{(q;q)_{r-i}}\frac{(q^{\mu_{1}};q)_{i}}{(q;q)_{i}}$
$\hskip 14.22636pt\cdot{}_{3}\phi_{2}\left[\begin{matrix}q^{\lambda_{1}-\alpha_{1}},q^{\gamma_{1}-\alpha_{1}},q^{i-r}\\ q^{\gamma_{1}+\lambda_{1}-\alpha_{1}-\mu_{1}},q^{1-r-\alpha_{1}}\end{matrix};q,q^{1-i-\mu_{1}}\right]q^{(r-i)\mu_{1}},$
$w_{2}(j,s;q)$ $:=\frac{(q^{\beta_{2}},q;q)_{s}}{(q^{\gamma_{2}},q^{\lambda_{2}};q)_{s}}\frac{(q^{\gamma_{2}+\lambda_{2}-\beta_{2}-\mu_{2}};q)_{s-j}}{(q;q)_{s-j}}\frac{(q^{\mu_{2}};q)_{j}}{(q;q)_{j}}$
$\hskip 14.22636pt\cdot{}_{3}\phi_{2}\left[\begin{matrix}q^{\lambda_{2}-\beta_{2}},q^{\gamma_{2}-\beta_{2}},q^{j-s}\\ q^{\gamma_{2}+\lambda_{2}-\beta_{2}-\mu_{2}},q^{1-s-\beta_{2}}\end{matrix};q,q^{1-j-\mu_{2}}\right]q^{(s-j)\mu_{2}}$
and
$w_{3}(k,t;q):=\frac{(q;q)_{t}}{(q^{\gamma_{3}};q)_{t}}\frac{(q^{\gamma_{3}-\mu_{3}};q)_{t-k}}{(q;q)_{t-k}}\frac{(q^{\mu_{3}};q)_{k}}{(q;q)_{k}}q^{(t-k)\mu_{3}}.$
Proof.
For convenience, we denote the right-hand side of (4.7) by $S$. By (2), it is not difficult to see that
$\phi^{(3)}\left[\begin{matrix}-::-;q^{\alpha_{2}};q^{\beta_{1}}:q^{\lambda_{1}},q^{-i};q^{\lambda_{2}},q^{-j};q^{-k}\\ -::-;-;-:q^{\mu_{1}},\delta_{1};q^{\mu_{2}},\delta_{2};q^{\mu_{3}},\delta_{3}\end{matrix};q,q,q,q\right]$
$\hskip 14.22636pt=\sum_{p=0}^{k}\frac{(q^{\alpha_{2}},q^{\beta_{1}},q^{-k};q)_{p}}{(q^{\mu_{3}},\delta_{3},q;q)_{p}}q^{p}{}_{3}\phi_{2}\left[\begin{matrix}q^{\beta_{1}+p},q^{\lambda_{1}},q^{-i}\\ q^{\mu_{1}},\delta_{1}\end{matrix};q,q\right]{}_{3}\phi_{2}\left[\begin{matrix}q^{\alpha_{2}+p},q^{\lambda_{2}},q^{-j}\\ q^{\mu_{2}},\delta_{2}\end{matrix};q,q\right].$
Then
$S$ $=\sum_{i=0}^{r}\sum_{j=0}^{s}\sum_{k=0}^{t}w_{1}(i,r;q)w_{2}(j,s;q)w_{3}(k,t;q)\sum_{p=0}^{k}\frac{(q^{\alpha_{2}},q^{\beta_{1}},q^{-k};q)_{p}}{(q^{\mu_{3}},\delta_{3},q;q)_{p}}q^{p}$
$\hskip 14.22636pt\cdot{}_{3}\phi_{2}\left[\begin{matrix}q^{\beta_{1}+p},q^{\lambda_{1}},q^{-i}\\ q^{\mu_{1}},\delta_{1}\end{matrix};q,q\right]{}_{3}\phi_{2}\left[\begin{matrix}q^{\alpha_{2}+p},q^{\lambda_{2}},q^{-j}\\ q^{\mu_{2}},\delta_{2}\end{matrix};q,q\right]$
$=\sum_{k=0}^{t}w_{3}(k,t;q)\sum_{p=0}^{k}\frac{(q^{\alpha_{2}},q^{\beta_{1}},q^{-k};q)_{p}}{(q^{\mu_{3}},\delta_{3},q;q)_{p}}q^{p}$
$\hskip 14.22636pt\sum_{i=0}^{r}w_{1}(i,r;q){}_{3}\phi_{2}\left[\begin{matrix}q^{\beta_{1}+p},q^{\lambda_{1}},q^{-i}\\ q^{\mu_{1}},\delta_{1}\end{matrix};q,q\right]$
$\hskip 14.22636pt\sum_{j=0}^{s}w_{2}(j,s;q){}_{3}\phi_{2}\left[\begin{matrix}q^{\alpha_{2}+p},q^{\lambda_{2}},q^{-j}\\ q^{\mu_{2}},\delta_{2}\end{matrix};q,q\right].$ (48)
Letting $\alpha\rightarrow q^{\alpha_{1}}$, $\beta\rightarrow q^{\beta_{1}+p}$, $\gamma\rightarrow q^{\gamma_{1}}$, $\delta\rightarrow\delta_{1}$, $\mu\rightarrow q^{\lambda_{1}}$, $\lambda\rightarrow q^{\alpha_{1}}$ and $\nu\rightarrow q^{\mu_{1}}$ in (2), we have
$\sum_{i=0}^{r}w_{1}(i,r;q){}_{3}\phi_{2}\left[\begin{matrix}q^{\beta_{1}+p},q^{\lambda_{1}},q^{-i}\\ q^{\mu_{1}},\delta_{1}\end{matrix};q,q\right]$ $={}_{3}\phi_{2}\left[\begin{matrix}q^{\alpha_{1}},q^{\beta_{1}+p},q^{-r}\\ q^{\gamma_{1}},\delta_{1}\end{matrix};q,q\right]$
$=\sum_{i=0}^{r}\frac{(q^{\alpha_{1}},q^{\beta_{1}},q^{-r};q)_{i}}{(q^{\gamma_{1}},\delta_{1},q;q)_{i}}\frac{(q^{\beta_{1}+i};q)_{p}}{(q^{\beta_{1}};q)_{p}}q^{i}.$ (49)
Letting $\alpha\rightarrow q^{\alpha_{2}+p}$, $\beta\rightarrow q^{\beta_{2}}$, $\gamma\rightarrow q^{\gamma_{2}}$, $\delta\rightarrow\delta_{2}$, $\mu\rightarrow q^{\lambda_{2}}$, $\lambda\rightarrow q^{\beta_{2}}$ and $\nu\rightarrow q^{\mu_{2}}$ in (2), we obtain
$\sum_{j=0}^{s}w_{2}(j,s;q){}_{3}\phi_{2}\left[\begin{matrix}q^{\alpha_{2}+p},q^{\lambda_{2}},q^{-j}\\ q^{\mu_{2}},\delta_{2}\end{matrix};q,q\right]$ $={}_{3}\phi_{2}\left[\begin{matrix}q^{\alpha_{2}+p},q^{\beta_{2}},q^{-s}\\ q^{\gamma_{2}},\delta_{2}\end{matrix};q,q\right]$
$=\sum_{j=0}^{s}\frac{(q^{\alpha_{2}},q^{\beta_{2}},q^{-s};q)_{j}}{(q^{\gamma_{2}},\delta_{2},q;q)_{j}}\frac{(q^{\alpha_{2}+j};q)_{p}}{(q^{\alpha_{2}};q)_{p}}q^{j}.$ (50)
Now, substituting (4.2) and (4.2) into (48) gives
$S$ $=\sum_{i=0}^{r}\frac{(q^{\alpha_{1}},q^{\beta_{1}},q^{-r};q)_{i}}{(q^{\gamma_{1}},\delta_{1},q;q)_{i}}q^{i}\sum_{j=0}^{s}\frac{(q^{\alpha_{2}},q^{\beta_{2}},q^{-s};q)_{j}}{(q^{\gamma_{2}},\delta_{2},q;q)_{j}}q^{j}$
$\hskip 14.22636pt\sum_{k=0}^{t}w_{3}(k,t;q)\sum_{p=0}^{k}\frac{(q^{\alpha_{2}+j},q^{\beta_{1}+i},q^{-k};q)_{p}}{(q^{\mu_{3}},\delta_{3},q;q)_{p}}q^{p}.$ (51)
Letting $\alpha\rightarrow q^{\alpha_{2}+j}$, $\beta\rightarrow q^{\beta_{1}+i}$, $\gamma\rightarrow q^{\gamma_{3}}$, $\delta\rightarrow\delta_{3}$, $\mu=\lambda$ and $\nu\rightarrow q^{\mu_{3}}$ in (2) gives
$\sum_{k=0}^{t}w_{3}(k,t;q)\sum_{p=0}^{k}\frac{(q^{\alpha_{2}+j},q^{\beta_{1}+i},q^{-k};q)_{p}}{(q^{\mu_{3}},\delta_{3},q;q)_{p}}q^{p}={}_{3}\phi_{2}\left[\begin{matrix}q^{\alpha_{2}+j},q^{\beta_{1}+i},q^{-n}\\ q^{\gamma_{3}},\delta_{3}\end{matrix};q,q\right].$ (52)
Substituting (52) into (4.2), we obtain
$S$ $=\sum_{i=0}^{r}\frac{(q^{\alpha_{1}},q^{\beta_{1}},q^{-r};q)_{i}}{(q^{\gamma_{1}},\delta_{1},q;q)_{i}}q^{i}\sum_{j=0}^{s}\frac{(q^{\alpha_{2}},q^{\beta_{2}},q^{-s};q)_{j}}{(q^{\gamma_{2}},\delta_{2},q;q)_{j}}q^{j}\sum_{k=0}^{t}\frac{(q^{\alpha_{2}+j},q^{\beta_{1}+i},q^{-t};q)_{k}}{(q^{\gamma_{3}},\delta_{3},q;q)_{k}}q^{k}$
$=\sum_{i=0}^{r}\sum_{j=0}^{s}\sum_{k=0}^{t}(q^{\alpha_{2}};q)_{j+k}(q^{\beta_{1}};q)_{k+i}\frac{(q^{\alpha_{1}},q^{-r};q)_{i}}{(q^{\gamma_{1}},\delta_{1},q;q)_{i}}\frac{(q^{\beta_{2}},q^{-s};q)_{j}}{(q^{\gamma_{2}},\delta_{2},q;q)_{j}}\frac{(q^{-t};q)_{k}}{(q^{\gamma_{3}},\delta_{3},q;q)_{k}}q^{i+j+k}$
Interpreting the above triple sum as the $\phi^{(3)}$-function leads us to the formula (4.7). ∎
By letting $i\rightarrow r-i$, $j\rightarrow s-j$, $k\rightarrow t-k$, $\delta_{1}\rightarrow q^{1-r}/x$, $\delta_{2}\rightarrow q^{1-s}/y$, $\delta_{3}\rightarrow q^{1-t}/z$ in (4.7), we have
$\phi^{(3)}\left[\begin{matrix}-::-;q^{\alpha_{2}};q^{\beta_{1}}:q^{\alpha_{1}},q^{-r};q^{\beta_{2}},q^{-s};q^{-t}\\ -::-;-;-:q^{\gamma_{1}},q^{1-r}/x;q^{\gamma_{2}},q^{1-s}/y;q^{\gamma_{3}},q^{1-t}/z\end{matrix};q,q,q,q\right]$
$=\sum_{i=0}^{r}\sum_{j=0}^{s}\sum_{k=0}^{t}w_{1}(r-i,r;q)w_{2}(s-j,s;q)w_{3}(t-k,t;q)$
$\hskip 14.22636pt\cdot\phi^{(3)}\left[\begin{matrix}-::-;q^{\alpha_{2}};q^{\beta_{1}}:q^{\lambda_{1}},q^{-i};q^{\lambda_{2}},q^{-j};q^{-k}\\ -::-;-;-:q^{\mu_{1}},q^{1-r}/x;q^{\mu_{2}},q^{1-s}/y;q^{\mu_{3}},q^{1-t}/z\end{matrix};q,q,q,q\right].$ (53)
Note that
$\lim_{r\rightarrow+\infty}w_{1}(r-i,r;q)$ $=\frac{(q^{\alpha_{1}},q^{\mu_{1}};q)_{\infty}}{(q^{\gamma_{1}},q^{\lambda_{1}};q)_{\infty}}\frac{(q^{\gamma_{1}+\lambda_{1}-\alpha_{1}-\mu_{1}};q)_{i}}{(q;q)_{i}}q^{i\mu_{1}}$
$\cdot{}_{3}\phi_{1}\left[\begin{matrix}q^{\lambda_{1}-\alpha_{1}},q^{\gamma_{1}-\alpha_{1}},q^{-i}\\ q^{\gamma_{1}+\lambda_{1}-\alpha_{1}-\mu_{1}}\end{matrix};q,q^{\alpha_{1}-\mu_{1}+i}\right],$ (54)
$\lim_{s\rightarrow+\infty}w_{2}(s-j,s;q)$ $=\frac{(q^{\beta_{2}},q^{\mu_{2}};q)_{\infty}}{(q^{\gamma_{2}},q^{\lambda_{2}};q)_{\infty}}\frac{(q^{\gamma_{2}+\lambda_{2}-\beta_{2}-\mu_{2}};q)_{j}}{(q;q)_{j}}q^{j\mu_{3}}$
$\cdot{}_{3}\phi_{1}\left[\begin{matrix}q^{\lambda_{2}-\beta_{2}},q^{\gamma_{2}-\beta_{2}},q^{-j}\\ q^{\gamma_{2}+\lambda_{2}-\beta_{2}-\mu_{2}}\end{matrix};q,q^{\beta_{2}-\mu_{2}+j}\right]$ (55)
and
$\lim_{t\rightarrow+\infty}w_{3}(t-k,t;q):=\frac{(q^{\mu_{3}};q)_{\infty}}{(q^{\gamma_{3}};q)_{\infty}}\frac{(q^{\gamma_{3}-\mu_{3}};q)_{k}}{(q;q)_{k}}q^{k\mu_{3}}.$ (56)
Letting $r,s,t\rightarrow+\infty$ in (4.2) and making use of (54), (55) and (56), we obtain
$\widetilde{\Phi}_{K}[\alpha_{1},\alpha_{2},\alpha_{2},\beta_{1},\beta_{2},\beta_{1};\gamma_{1},\gamma_{2},\gamma_{3};q,x,y,z]$
$=\frac{(q^{\alpha_{1}},q^{\mu_{1}},q^{\gamma_{1}-\lambda_{1}-\alpha_{1}-\mu_{1}};q)_{\infty}}{(q^{\gamma_{1}},q^{\lambda_{1}},q;q)_{\infty}}\frac{(q^{\beta_{2}},q^{\mu_{2}},q^{\gamma_{2}+\lambda_{2}-\beta_{2}-\mu_{2}};q)_{\infty}}{(q^{\gamma_{2}},q^{\lambda_{2}},q;q)_{\infty}}\frac{(q^{\mu_{3}},q^{\gamma_{3}-\mu_{3}};q)_{\infty}}{(q^{\gamma_{3}},q;q)_{\infty}}$
$\hskip 14.22636pt\cdot\sum_{i,j,k=0}^{\infty}\frac{(q^{1+i};q)_{\infty}}{(q^{\gamma_{1}+\lambda_{1}-\alpha_{1}-\mu_{1}+i};q)_{\infty}}\frac{(q^{1+j};q)_{\infty}}{(q^{\gamma_{2}+\lambda_{2}-\beta_{2}-\mu_{2}+j};q)_{\infty}}\frac{(q^{1+k};q)_{\infty}}{(q^{\gamma_{3}-\mu_{3}+k};q)_{\infty}}q^{i\mu_{1}+j\mu_{2}+k\mu_{3}}$
$\hskip 14.22636pt\cdot{}_{3}\phi_{1}\left[\begin{matrix}q^{\lambda_{1}-\alpha_{1}},q^{\gamma_{1}-\alpha_{1}},q^{-i}\\ q^{\gamma_{1}+\lambda_{1}-\alpha_{1}-\mu_{1}}\end{matrix};q,q^{\alpha_{1}-\mu_{1}+i}\right]{}_{3}\phi_{1}\left[\begin{matrix}q^{\lambda_{2}-\beta_{2}},q^{\gamma_{2}-\beta_{2}},q^{-j}\\ q^{\gamma_{2}+\lambda_{2}-\beta_{2}-\mu_{2}}\end{matrix};q,q^{\beta_{2}-\mu_{2}+j}\right]$
$\hskip 14.22636pt\cdot\widetilde{\Phi}_{K}[\lambda_{1},\alpha_{2},\alpha_{2},\beta_{1},\lambda_{2},\beta_{1};\mu_{1},\mu_{2},\mu_{3};q,xq^{i},yq^{j},zq^{k}].$
Using definition (12) of the multiple $q$-integral, the expression can be further reduced to (4.5).
4.3 A $q$-analogue of Theorem 1.1
Let $f:\mathbb{C}\rightarrow\mathbb{C}$ be a function of one variable. We define the $q$-shift operator $\mathbb{E}_{q}$ by $\mathbb{E}_{q}f(x)=f(qx)$ $(0<q<1)$ [29, Definition 1.3]. Inductively, for any $n\in\mathbb{Z}_{\geq 1}$,
$\mathbb{E}_{q}^{n}f(x)=f(q^{n}x)$
and in particular, $\mathbb{E}_{q}^{0}f(x)=f(x)$. In addition, $\lim_{q\rightarrow 1^{-}}\mathbb{E}_{q}^{n}f(x)=f(x)$ . For $k$-dimensional case, we use the notation $\mathbb{E}_{q,x_{i}}$ to denote the operator acting on the $i$-th variable of $f$, namely,
$\mathbb{E}_{q,x_{i}}f(x_{1},\cdots,x_{i},\cdots,x_{k})=f(x_{1},\cdots,qx_{i},\cdots,x_{k}),\penalty 10000\ i=1,\cdots,k.$
We shall frequently use the formula
$\mathbb{E}_{q,x}^{n}\mathbb{E}_{q,y}^{m}f(x,y)=f(q^{n}x,q^{m}y),\penalty 10000\ \penalty 10000\ \penalty 10000\ (n,m)\in\mathbb{Z}_{\geq 0}^{2}.$ (57)
If $\mathcal{E}(x)=\sum_{n=0}^{\infty}a_{n}x^{n}$ $(|x|<R)$ is an analytic functions, then
$\mathcal{E}(t\penalty 10000\ \mathbb{E}_{q,x})f(x)=\sum_{n=0}^{\infty}a_{n}t^{n}f(q^{n}x),\penalty 10000\ \penalty 10000\ \penalty 10000\ |t|<R,$
where $f(x)$ is also assumed to be bounded (or, in many situations, analytic) in a neighborhood of 0. We have
$\lim_{q\rightarrow 1^{-}}\mathcal{E}(t\penalty 10000\ \mathbb{E}_{q,x})f(x)=\mathcal{E}(t)f(x).$
The limit helps overcome the difficulty in deriving the $q$-analogue of Theorem 1.1.
Theorem 4.8.
Let $\Re(\alpha_{1}+\eta_{1})>\Re(\lambda_{1})>0$, $\Re(\beta_{2}+\mu_{2})>\Re(\lambda_{2})>0$ and $\Re(\gamma_{3})>\Re(\beta_{1})>0$. Then we have
$\widetilde{\Phi}_{K}[\alpha_{1},\alpha_{2},\alpha_{2},\beta_{1},\beta_{2},\beta_{1};\alpha_{1}+\eta_{1},\beta_{2}+\mu_{2},\gamma_{3};q,x,y,z]=\int_{[0,1]^{3}}\frac{(uxq^{\lambda_{3}},vyq^{\eta_{2}};q)_{\infty}}{(ux,vy;q)_{\infty}}$
$\hskip 14.22636pt\cdot\phi^{(3)}\Bigg[\begin{matrix}-::-;q^{\eta_{2}};q^{\lambda_{3}}:q^{\lambda_{1}-\eta_{1}},u^{-1};q^{\lambda_{2}-\mu_{2}},v^{-1};-\\ -::-;-;-:q^{\lambda_{1}},q/(ux);q^{\lambda_{2}},q/(vy);uxq^{\lambda_{3}},vyq^{\eta_{2}},q^{\lambda_{3}}\end{matrix};$
$\hskip 42.67912ptq,q,q,wzq^{\alpha_{2}-\eta_{2}}\mathbb{E}_{q,x}\mathbb{E}_{q,y}\Bigg]$
$\hskip 14.22636pt\cdot\widetilde{\Phi}_{K}\big[\alpha_{1},\alpha_{2}-\eta_{2},\alpha_{2}-\eta_{2},\beta_{1}-\lambda_{3},\beta_{2},\beta_{1}-\lambda_{3};\alpha_{1}-\lambda_{1}+\eta_{1},\beta_{2}-\lambda_{2}+\mu_{2},\beta_{1}-\lambda_{3};$
$\hskip 42.67912ptq,uxq^{\lambda_{3}},vyq^{\eta_{2}},wz\big]\mathrm{d}\mu_{\alpha_{1}-\lambda_{1}+\eta_{1},\lambda_{1}}(u;q)\mathrm{d}\mu_{\beta_{2}-\lambda_{2}+\mu_{2},\lambda_{2}}(v;q)\mathrm{d}\mu_{\beta_{1},\gamma_{3}-\beta_{1}}(w;q).$ (58)
Proof.
We use the method from Section 3.1. Let II denote the triple $q$-integral in (4.8). By the definition (2) of the $\phi^{(3)}$-function, we have
$\phi^{(3)}\left[\begin{matrix}-::-;q^{\eta_{2}};q^{\lambda_{3}}:q^{\lambda_{1}-\eta_{1}},u^{-1};q^{\lambda_{2}-\mu_{2}},v^{-1};-\\ -::-;-;-:q^{\lambda_{1}},q/(ux);q^{\lambda_{2}},q/(vy);uxq^{\lambda_{3}},vyq^{\eta_{2}},q^{\lambda_{3}}\end{matrix};q,q,q,wzq^{\alpha_{2}-\eta_{2}}\mathbb{E}_{q,x}\mathbb{E}_{q,y}\right]$
$\hskip 14.22636pt=\sum_{k=0}^{\infty}\frac{(q^{\eta_{2}};q)_{k}}{(uxq^{\lambda_{3}},vyq^{\eta_{2}};q)_{k}}{}_{3}\phi_{2}\left[\begin{matrix}q^{\lambda_{3}+k},q^{\lambda_{1}-\eta_{1}},u^{-1}\\ q^{\lambda_{1}},q/(ux)\end{matrix};q,q\right]$
$\hskip 42.67912pt\cdot{}_{3}\phi_{2}\left[\begin{matrix}q^{\eta_{2}+k},q^{\lambda_{2}-\mu_{2}},v^{-1}\\ q^{\lambda_{2}},q/(vy)\end{matrix};q,q\right]\frac{(wzq^{\alpha_{2}-\eta_{2}}\mathbb{E}_{q,x}\mathbb{E}_{q,y})^{k}}{(q;q)_{k}}.$
Then, in view of (57) and (14), we have
$I$ $=\int_{[0,1]^{3}}\frac{(uxq^{\lambda_{3}},vyq^{\eta_{2}};q)_{\infty}}{(ux,vy;q)_{\infty}}\sum_{k=0}^{\infty}\frac{(q^{\eta_{2}};q)_{k}}{(uxq^{\lambda_{3}},vyq^{\eta_{2}};q)_{k}}$
$\hskip 14.22636pt\cdot{}_{3}\phi_{2}\left[\begin{matrix}q^{\lambda_{3}+k},q^{\lambda_{1}-\eta_{1}},u^{-1}\\ q^{\lambda_{1}},q/(ux)\end{matrix};q,q\right]{}_{3}\phi_{2}\left[\begin{matrix}q^{\eta_{2}+k},q^{\lambda_{2}-\mu_{2}},v^{-1}\\ q^{\lambda_{2}},q/(vy)\end{matrix};q,q\right]\frac{(wzq^{\alpha_{2}-\eta_{2}})^{k}}{(q;q)_{k}}$
$\hskip 14.22636pt\cdot\widetilde{\Phi}_{K}\big[\alpha_{1},\alpha_{2}-\eta_{2},\alpha_{2}-\eta_{2},\beta_{1}-\lambda_{3},\beta_{2},\beta_{1}-\lambda_{3};\alpha_{1}-\lambda_{1}+\eta_{1},\beta_{2}-\lambda_{2}+\mu_{2},\beta_{1}-\lambda_{3};$
$\hskip 28.45274ptuxq^{k+\lambda_{3}},vyq^{k+\eta_{2}},wz\big]\mathrm{d}\mu_{\alpha_{1}-\lambda_{1}+\eta_{1},\lambda_{1}}(u;q)\mathrm{d}\mu_{\beta_{2}-\lambda_{2}+\mu_{2},\lambda_{2}}(v;q)\mathrm{d}\mu_{\beta_{1},\gamma_{3}-\beta_{1}}(w;q)$
$=\int_{[0,1]^{3}}\sum_{k=0}^{\infty}\frac{(q^{\eta_{2}};q)_{k}}{(q;q)_{k}}\frac{(uxq^{k+\lambda_{3}},vyq^{k+\eta_{2}};q)_{\infty}}{(ux,vy;q)_{\infty}}(wzq^{\alpha_{2}-\eta_{2}})^{k}$
$\hskip 14.22636pt\cdot{}_{3}\phi_{2}\left[\begin{matrix}q^{\lambda_{3}+k},q^{\lambda_{1}-\eta_{1}},u^{-1}\\ q^{\lambda_{1}},q/(ux)\end{matrix};q,q\right]{}_{3}\phi_{2}\left[\begin{matrix}q^{\eta_{2}+k},q^{\lambda_{2}-\mu_{2}},v^{-1}\\ q^{\lambda_{2}},q/(vy)\end{matrix};q,q\right]$
$\hskip 14.22636pt\cdot\sum_{\ell=0}^{\infty}\frac{(q^{\alpha_{2}-\eta_{2}};q)_{\ell}}{(q;q)_{\ell}}{}_{2}\widetilde{\phi}_{1}\left[\begin{matrix}\beta_{1}-\lambda_{3}+\ell,\alpha_{1}\\ \alpha_{1}-\lambda_{1}+\eta_{1}\end{matrix};q,uxq^{k+\lambda_{3}}\right]{}_{2}\widetilde{\phi}_{1}\left[\begin{matrix}\alpha_{2}-\eta_{2}+\ell,\beta_{2}\\ \beta_{2}-\lambda_{2}+\mu_{2}\end{matrix};q,vyq^{k+\eta_{2}}\right]$
$\hskip 14.22636pt\cdot(wz)^{\ell}\mathrm{d}\mu_{\alpha_{1}-\lambda_{1}+\eta_{1},\lambda_{1}}(u;q)\mathrm{d}\mu_{\beta_{2}-\lambda_{2}+\mu_{2},\lambda_{2}}(v;q)\mathrm{d}\mu_{\beta_{1},\gamma_{3}-\beta_{1}}(w;q)$
$=\sum_{k=0}^{\infty}\frac{(q^{\eta_{2}};q)_{k}}{(q;q)_{k}}(zq^{\alpha_{2}-\eta_{2}})^{k}\sum_{\ell=0}^{\infty}\frac{(q^{\alpha_{2}-\eta_{2}};q)_{\ell}}{(q;q)_{\ell}}z^{\ell}\int_{0}^{1}w^{k+\ell}\mathrm{d}\mu_{\beta_{1},\gamma_{3}-\beta_{1}}(w;q)$
$\hskip 14.22636pt\cdot\int_{0}^{1}\frac{(uxq^{\lambda_{3}+k};q)_{\infty}}{(ux;q)_{\infty}}{}_{2}\widetilde{\phi}_{1}\left[\begin{matrix}\beta_{1}-\lambda_{3}+\ell,\alpha_{1}\\ \alpha_{1}-\lambda_{1}+\eta_{1}\end{matrix};q,uxq^{\lambda_{3}+k}\right]$
$\hskip 56.9055pt\cdot{}_{3}\phi_{2}\left[\begin{matrix}q^{\lambda_{3}+k},q^{\lambda_{1}-\eta_{1}},u^{-1}\\ q^{\lambda_{1}},q/(ux)\end{matrix};q,q\right]\mathrm{d}\mu_{\alpha_{1}-\lambda_{1}+\eta_{1},\lambda_{1}}(u;q)$
$\hskip 14.22636pt\cdot\int_{0}^{1}\frac{(vyq^{\eta_{2}+k};q)_{\infty}}{(vy;q)_{\infty}}{}_{2}\widetilde{\phi}_{1}\left[\begin{matrix}\alpha_{2}-\eta_{2}+\ell,\beta_{2}\\ \beta_{2}-\lambda_{2}+\mu_{2}\end{matrix};q,vyq^{\eta_{2}+k}\right]$
$\hskip 56.9055pt\cdot{}_{3}\phi_{2}\left[\begin{matrix}q^{\eta_{2}+k},q^{\lambda_{2}-\mu_{2}},v^{-1}\\ q^{\lambda_{2}},q/(vy)\end{matrix};q,q\right]\mathrm{d}\mu_{\beta_{2}-\lambda_{2}+\mu_{2},\lambda_{2}}(v;q).$ (59)
By letting $\alpha\rightarrow\beta_{1}+k+\ell$, $\beta\rightarrow\alpha_{1}$, $\gamma\rightarrow\alpha_{1}+\eta_{1}$, $\lambda\rightarrow\alpha_{1}-\lambda_{1}+\eta_{1}$ and $\alpha^{\prime}\rightarrow\lambda_{3}+k$ in (2), we have
$\int_{0}^{1}\frac{(uxq^{\lambda_{3}+k};q)_{\infty}}{(ux;q)_{\infty}}{}_{2}\widetilde{\phi}_{1}\left[\begin{matrix}\beta_{1}-\lambda_{3}+\ell,\alpha_{1}\\ \alpha_{1}-\lambda_{1}+\eta_{1}\end{matrix};q,uxq^{\lambda_{3}+k}\right]$
$\hskip 14.22636pt\cdot{}_{3}\phi_{2}\left[\begin{matrix}q^{\lambda_{3}+k},q^{\lambda_{1}-\eta_{1}},u^{-1}\\ q^{\lambda_{1}},q/(ux)\end{matrix};q,q\right]\mathrm{d}\mu_{\alpha_{1}-\lambda_{1}+\eta_{1},\lambda_{1}}(u;q)={}_{2}\widetilde{\phi}_{1}\left[\begin{matrix}\beta_{1}+k+\ell,\alpha_{1}\\ \alpha_{1}+\eta_{1}\end{matrix};q,x\right],$ (60)
where $\Re(\alpha_{1}+\eta_{1})>\Re(\lambda_{1})>0$. Similarly, let $\alpha\rightarrow\alpha_{2}+k+\ell$, $\beta\rightarrow\beta_{2}$, $\gamma\rightarrow\beta_{2}+\mu_{2}$, $\lambda=\beta_{2}-\lambda_{2}+\mu_{2}$ and $\alpha^{\prime}\rightarrow\eta_{2}+k$ in (2), then
$\int_{0}^{1}\frac{(vyq^{\eta_{2}+k};q)_{\infty}}{(vy;q)_{\infty}}{}_{2}\widetilde{\phi}_{1}\left[\begin{matrix}\alpha_{2}-\eta_{2}+\ell,\beta_{2}\\ \beta_{2}-\lambda_{2}+\mu_{2}\end{matrix};q,vyq^{\eta_{2}+k}\right]$
$\hskip 14.22636pt\cdot{}_{3}\phi_{2}\left[\begin{matrix}q^{\eta_{2}+k},q^{\lambda_{2}-\mu_{2}},v^{-1}\\ q^{\lambda_{2}},q/(vy)\end{matrix};q,q\right]\mathrm{d}\mu_{\beta_{2}-\lambda_{2}+\mu_{2},\lambda_{2}}(v;q)={}_{2}\widetilde{\phi}_{1}\left[\begin{matrix}\alpha_{2}+k+\ell,\beta_{2}\\ \beta_{2}+\mu_{2}\end{matrix};q,y\right],$ (61)
where $\Re(\beta_{2}+\mu_{2})>\Re(\lambda_{2})>0$. Substituting (4.3) and (4.3) in (59) gives
$I$ $=\sum_{k=0}^{\infty}\frac{(q^{\eta_{2}};q)_{k}}{(q;q)_{k}}(q^{\alpha_{2}-\eta_{2}})^{k}\sum_{\ell=0}^{\infty}\frac{(q^{\alpha_{2}-\eta_{2}};q)_{\ell}}{(q;q)_{\ell}}\frac{(q^{\beta_{1}};q)_{k+\ell}}{(q^{\gamma_{3}};q)_{k+\ell}}z^{k+\ell}$
$\hskip 28.45274pt\cdot{}_{2}\widetilde{\phi}_{1}\left[\begin{matrix}\beta_{1}+k+\ell,\alpha_{1}\\ \alpha_{1}+\eta_{1}\end{matrix};q,x\right]{}_{2}\widetilde{\phi}_{1}\left[\begin{matrix}\alpha_{2}+k+\ell,\beta_{2}\\ \beta_{2}+\mu_{2}\end{matrix};q,y\right]$
$=\sum_{k=0}^{\infty}\frac{(q^{\beta_{1}};q)_{k}}{(q^{\gamma_{3}};q)_{k}}z^{k}{}_{2}\widetilde{\phi}_{1}\left[\begin{matrix}\beta_{1}+k,\alpha_{1}\\ \alpha_{1}+\eta_{1}\end{matrix};q,x\right]{}_{2}\widetilde{\phi}_{1}\left[\begin{matrix}\alpha_{2}+k,\beta_{2}\\ \beta_{2}+\mu_{2}\end{matrix};q,y\right]$
$\hskip 28.45274pt\cdot\sum_{\ell=0}^{k}\frac{(q^{\eta_{2}};q)_{k-\ell}}{(q;q)_{k-\ell}}\frac{(q^{\alpha_{2}-\eta_{2}};q)_{\ell}}{(q;q)_{\ell}}(q^{\alpha_{2}-\eta_{2}})^{k-\ell}$
$=\sum_{k=0}^{\infty}\frac{(q^{\alpha_{2}},q^{\beta_{1}};q)_{k}}{(q^{\gamma_{3}},q;q)_{k}}z^{k}{}_{2}\widetilde{\phi}_{1}\left[\begin{matrix}\beta_{1}+k,\alpha_{1}\\ \alpha_{1}+\eta_{1}\end{matrix};q,x\right]{}_{2}\widetilde{\phi}_{1}\left[\begin{matrix}\alpha_{2}+k,\beta_{2}\\ \beta_{2}+\mu_{2}\end{matrix};q,y\right].$ (62)
By interpreting the last series in (4.3) as the $q$-Saran function $\widetilde{\Phi}_{K}$, we obtain the desired result (4.8). ∎
Remark 4.9.
The formula (4.8) obtained here involves the $q$-shift operators $\mathbb{E}_{q,x}$ and $\mathbb{E}_{q,y}$, which makes it appears slightly different from the familiar integrals associated with $_{2}\phi_{1}$. It would therefore be interesting to explore whether alternative $q$-Erdélyi-type integrals for $\Phi_{K}$ can be derived through a different approach. We conjecture that the use of $q$-fractional integration by parts [2, p. 159, Theorem 5.14] may offer a viable path for addressing this problem.
Finally, we show that in some cases, the integrand in (4.8) can be simplified so that the operators $\mathbb{E}_{q,x}$ and $\mathbb{E}_{q,y}$ do not appear any more.
Corollary 4.10.
When $\min\{\Re(\alpha_{1}),\Re(\eta_{1}),\Re(\beta_{2}),\Re(\mu_{2})\}>0$ and $\Re(\gamma_{3})>\Re(\beta_{1})>0$, we have
$\widetilde{\Phi}_{K}[\alpha_{1},\alpha_{2},\alpha_{2},\beta_{1},\beta_{2},\beta_{1};\alpha_{1}+\eta_{1},\beta_{2}+\mu_{2},\gamma_{3};q,x,y,z]=\int_{[0,1]^{3}}\frac{(uxq^{\beta_{1}},vyq^{\alpha_{2}};q)_{\infty}}{(ux,vy;q)_{\infty}}$
$\hskip 28.45274pt\cdot{}_{3}\phi_{2}\left[\begin{matrix}q^{\alpha_{2}},0,0\\ uxq^{\beta_{1}},vyq^{\alpha_{2}}\end{matrix};q,wz\right]\mathrm{d}\mu_{\alpha_{1},\eta_{1}}(u;q)\mathrm{d}\mu_{\beta_{2},\mu_{2}}(v;q)\mathrm{d}\mu_{\beta_{1},\gamma_{3}-\beta_{1}}(w;q).$ (63)
Proof.
By letting $\lambda_{1}\rightarrow\eta_{1}$ and $\lambda_{2}\rightarrow\mu_{2}$ in (4.8), we obtain
$\widetilde{\Phi}_{K}[\alpha_{1},\alpha_{2},\alpha_{2},\beta_{1},\beta_{2},\beta_{1};\alpha_{1}+\eta_{1},\beta_{2}+\mu_{2},\gamma_{3};q,x,y,z]=\int_{[0,1]^{3}}\frac{(uxq^{\lambda_{3}},vyq^{\eta_{2}};q)_{\infty}}{(ux,vy;q)_{\infty}}$
$\hskip 14.22636pt\cdot\mathcal{K}(q;x,y,z;u,v,w)\mathrm{d}\mu_{\alpha_{1},\eta_{1}}(u;q)\mathrm{d}\mu_{\beta_{2},\mu_{2}}(v;q)\mathrm{d}\mu_{\beta_{1},\gamma_{3}-\beta_{1}}(w;q),$ (64)
where
$\mathcal{K}(q;x,y,z;u,v,w)$ $:={}_{3}\phi_{2}\left[\begin{matrix}q^{\eta_{2}},0,0\\ uxq^{\lambda_{3}},vyq^{\eta_{2}}\end{matrix};q,wzq^{\alpha_{2}-\eta_{2}}\mathbb{E}_{q,x}\mathbb{E}_{q,y}\right]$
$\hskip 28.45274pt\frac{(uxq^{\beta_{1}},vyq^{\alpha_{2}};q)_{\infty}}{(uxq^{\lambda_{3}},vyq^{\eta_{2}};q)_{\infty}}{}_{3}\phi_{2}\left[\begin{matrix}q^{\alpha_{2}-\eta_{2}},0,0\\ uxq^{\beta_{1}},vyq^{\alpha_{2}}\end{matrix};q,wz\right].$
To evaluate $\mathcal{K}(q;x,y,z;u,v,w)$, we first note that
$\mathcal{K}(q;x,y,z;u,v,w)$ $=\sum_{k=0}^{\infty}\frac{(q^{\eta_{2}};q)_{k}(wzq^{\alpha_{2}-\eta_{2}})^{k}}{(uxq^{\lambda_{3}},vyq^{\eta_{2}},q;q)_{k}}\frac{(uxq^{\beta_{1}+k},vyq^{\alpha_{2}+k};q)_{\infty}}{(uxq^{\lambda_{3}+k},vyq^{\eta_{2}+k};q)_{\infty}}$
$\hskip 14.22636pt\cdot\sum_{p=0}^{\infty}\frac{(q^{\alpha_{2}-\eta_{2}};q)_{p}(wz)^{p}}{(uxq^{\beta_{1}+k},vyq^{\alpha_{2}+k},q;q)_{p}}$
$=\frac{(uxq^{\beta_{1}},vyq^{\alpha_{2}};q)_{\infty}}{(uxq^{\lambda_{3}},vyq^{\eta_{2}};q)_{\infty}}\sum_{k=0}^{\infty}\sum_{p=0}^{\infty}\frac{(q^{\eta_{2}};q)_{k}(q^{\alpha_{2}-\eta_{2}};q)_{p}q^{k(\alpha_{2}-\eta_{2})}(wz)^{k+p}}{(uxq^{\beta_{1}},vyq^{\alpha_{2}};q)_{k+p}(q;q)_{k}(q;q)_{p}}$
$=\frac{(uxq^{\beta_{1}},vyq^{\alpha_{2}};q)_{\infty}}{(uxq^{\lambda_{3}},vyq^{\eta_{2}};q)_{\infty}}\sum_{k=0}^{\infty}\frac{(wz)^{k}}{(uxq^{\beta_{1}},vyq^{\alpha_{2}},q;q)_{k}}$
$\hskip 14.22636pt\cdot\sum_{p=0}^{k}\left[\begin{matrix}k\\ p\end{matrix}\right]_{q}(q^{\eta_{2}};q)_{k-p}(q^{\alpha_{2}-\eta_{2}};q)_{p}q^{(k-p)(\alpha_{2}-\eta_{2})},$
where $\left[\begin{matrix}k\\ p\end{matrix}\right]_{q}$ denotes the $q$-binomial coefficient [19, p. 24]. Then the use of the $q$-binomial theorem [19, p. 25] gives
$\mathcal{K}(q;x,y,z;u,v,w)$ $=\frac{(uxq^{\beta_{1}},vyq^{\alpha_{2}};q)_{\infty}}{(uxq^{\lambda_{3}},vyq^{\eta_{2}};q)_{\infty}}\sum_{k=0}^{\infty}\frac{(q^{\alpha_{2}};q)_{k}(wz)^{k}}{(uxq^{\beta_{1}},vyq^{\alpha_{2}},q;q)_{k}}$
$=\frac{(uxq^{\beta_{1}},vyq^{\alpha_{2}};q)_{\infty}}{(uxq^{\lambda_{3}},vyq^{\eta_{2}};q)_{\infty}}{}_{3}\phi_{2}\left[\begin{matrix}q^{\alpha_{2}},0,0\\ uxq^{\beta_{1}},vyq^{\alpha_{2}}\end{matrix};q,wz\right].$ (65)
Substituting (4.3) into (4.3) we obtain (4.10). ∎
5 Appendix
Type of function Reference
11 variable $_{2}F_{1}$ Erdélyi (1939) [10], [12]
Joshi-Vyas (2003) [21]
$_{p}F_{q}$ Joshi-Vyas (2003) [21]
Luo-Raina (2017) [30]
$q$-analogues Gasper (2000) [18]
Joshi-Vyas (2006) [22]
Vyas (2024) [49]
discrete analogues Vyas-Bhatnagar-Fatawat-Suthar-Purohit (2022) [50]
Bhatnagar-Vyas (2022) [3]
others Laine (1982) [28]
$n$ variables Appell $F_{1}$ Manocha (1967) [34]
($n\geq 2$) Mittal (1977) [37]
$F_{2}$ Koschmieder (1947) [26]
Manocha (1967) [34]
Mittal (1977) [37]
$F_{3}$ Mittal (1977) [37]
$F_{4}$ Manocha (1965) [35]
Mittal (1977) [37]
Lauricella $F_{A}$ Manocha-Sharma (1969) [36]
Chandel (1971) [7]
$F_{B}$ Chandel (1971) [7]
$F_{C}$ Chandel (1971) [7]
$F_{D}$ Koschmieder (1962) [27]
Manocha-Sharma (1969) [36]
Chandel (1971) [7]
Khudozhnikov (2003) [24]
Saran $F_{M}$ Manocha-Sharma (1969) [36]
$F_{K}$ Luo-Raina (2021) [31]
Luo-Xu-Raina (2022) [32]
$q$-analogues the present paper
Funding
No funding was received for conducting this study.
Conflicts of interest
The authors declare that there is no conflict of interest.
Data Availability
This manuscript has no associated data.
Acknowledgements
We are grateful to Peng-Cheng Hang for his careful reading of the initial manuscript, for correcting several typos, and for providing valuable suggestions.

### v2

@@
 Abstract
-In this paper, we revisit the recent result of Luo, Xu, and Raina [Fractal Frac.
+In this paper, we revisit the recent result of Luo, Xu, and Raina [Fractal Fract.
 6 (3) (2022)] on an Erdélyi-type integral for Saran’s three-variable hypergeometric function $F_{K}$.
@@
 People’s Republic of China.
-E-mail: mathwinnie@live.com, mathwinnie@dhu.edu.cn
+E-mail: guoliangjia77@outlook.com
 2School of Mathematics and Statistics,
@@
 People’s Republic of China.
-E-mail: guoliangjia77@outlook.com
+E-mail: mathwinnie@live.com, mathwinnie@dhu.edu.cn
 1 Introduction
@@
 $_{2}F_{1}\left[\begin{matrix}\alpha,\beta\\ \gamma\end{matrix};z\right]:=\sum_{n=0}^{\infty}\frac{(\alpha)_{n}(\beta)_{n}}{(\gamma)_{n}}\frac{z^{n}}{n!}\penalty 10000\ \penalty 10000\ \penalty 10000\ (|z|<1),$
-where $\alpha,\beta\in\mathbb{C}$ and $\gamma\notin\mathbb{Z}_{\leq 0}:=\{0,-1,-2,\cdots\}$ (see [13, p. 56] and [38, p. 384]). It was shown by Euler that [13, p. 59, Eq. (10)]
+where $\alpha,\beta\in\mathbb{C}$ and $\gamma\notin\mathbb{Z}_{\leq 0}:=\{0,-1,-2,\cdots\}$ (see [14, p. 56] and [39, p. 384]). It was shown by Euler that [14, p. 59, Eq. (10)]
 ${}_{2}F_{1}\left[\begin{matrix}\alpha,\beta\\ \gamma\end{matrix};z\right]=$ $\int_{0}^{1}(1-zt)^{-\alpha}\mathrm{d}\mu_{\beta,\gamma-\beta}(t)\penalty 10000\ \penalty 10000\ \penalty 10000\ (\Re(\gamma)>\Re(\beta)>0),$ (1)
 ${}_{2}F_{1}\left[\begin{matrix}\alpha,\beta\\ \gamma\end{matrix};z\right]=$ $\int_{0}^{1}(1-zt)^{-\beta}\mathrm{d}\mu_{\alpha,\gamma-\alpha}(t)\penalty 10000\ \penalty 10000\ \penalty 10000\ (\Re(\gamma)>\Re(\alpha)>0).$ (2)
-Here and throughout the paper, we let $\mu_{\alpha,\beta}$ be the Dirichlet measure defined by [6, p. 64, Definition 4.4-1]
+Here and throughout the paper, we let $\mu_{\alpha,\beta}$ be the Dirichlet measure defined by [7, p. 64, Definition 4.4-1]
 $\mathrm{d}\mu_{\alpha,\beta}(t):=\mathfrak{m}_{\alpha,\beta}(t)\mathrm{d}t,$ (3)
@@
 We are particularly interested in Erdélyi’s and Poole’s methods.
-Erdélyi’s method (see [11, p. 272–273]) is to substitute
+Erdélyi’s method (see [12, p. 272–273]) is to substitute
 $(1-zt)^{-\alpha}=\int_{0}^{1}(1-zts)^{-\gamma}\mathrm{d}\mu_{\alpha,\gamma-\alpha}(s)\penalty 10000\ \penalty 10000\ \penalty 10000\ (\Re(\gamma)>\Re(\alpha)>0)$
 into the first integral (1) to obtain the double integral
-$\int_{0}^{1}\int_{0}^{1}(1-zts)^{-\gamma}\mathrm{d}\mu_{\beta,\gamma-\beta}(t)\mathrm{d}\mu_{\alpha,\gamma-\alpha}(s),$
-which is symmetric in $\alpha$ and $\beta$. If we integrate with respect to $t$ first, then we obtain the second integral (2).
+$\int_{0}^{1}\int_{0}^{1}(1-zts)^{-\gamma}\mathrm{d}\mu_{\beta,\gamma-\beta}(t)\mathrm{d}\mu_{\alpha,\gamma-\alpha}(s),$ (4)
+which is symmetric in $\alpha$ and $\beta$.
+If we integrate with respect to $t$ first, then we obtain the second integral (2).
+Note that after introducing the new variables $x_{1}=s/(1-s)$ and $x_{2}=t/(1-t)$ and making the substitution $z\rightarrow 1-z$ in Erdélyi’s formula (4), we obtain the double integral
+$\int_{0}^{\infty}\int_{0}^{\infty}\frac{x_{1}^{\alpha}x_{2}^{\beta}}{(1+x_{1}+x_{2}+zx_{1}x_{2})^{\gamma}}\frac{\mathrm{d}x_{1}\mathrm{d}x_{2}}{x_{1}x_{2}}=\frac{\Gamma(\beta)\Gamma(\gamma-\beta)\Gamma(\alpha)\Gamma(\gamma-\alpha)}{(\Gamma(\gamma))^{2}}{}_{2}F_{1}\left[\begin{matrix}\alpha,\beta\\ \gamma\end{matrix};1-z\right].$ (5)
+The integral on the left-hand side of (5) is also known as the Euler-Mellin integral (see [3] and [50]).
 Poole’s method [6] is to apply the integration by parts to the contour integral representation of $_{2}F_{1}$, which is very delightful.
@@
 $\hskip 28.45274pt(\min\{\Re(\nu),\Re(\lambda),\Re(\gamma-\lambda+\eta-\nu)\}>0).$ (9)
-In (7), $\mu_{\alpha,\beta,\gamma,\eta}$ is the hypergeometric measure defined by [32, p. 6, Eq. (11)]
+In (9), $\mu_{\alpha,\beta,\gamma,\eta}$ is the hypergeometric measure defined by [33, p. 6, Eq. (11)]
 $\mathrm{d}\mu_{\alpha,\beta,\gamma,\eta}(t):=\mathfrak{m}_{\alpha,\beta,\gamma,\eta}(t)\mathrm{d}t,$ (10)
@@
 Gasper [18] has shown that (8) provides a natural way to the formulas of Dirichlet-Mehler type for the Jacobi polynomials and the generalized Legendre functions.
+Virchenko and Fedotova also frequently used the integrals (7) and (8) in their study of the generalized Legendre functions (see [51], [52] and [53]).
 Sprinkhuizen-Kuyper [48] used (8) to prove a compositional property for her fractional integral operator $I_{\nu}^{\mu,\lambda}$.
@@
 We refer to integrals that have a form similar to integrals (7)–(9) as Erdélyi-type integrals.
-More precisely, for a given hypergeometric function, its Erdélyi-type integrals can generally express it as an integral of the product of two functions of the same type (see [32, p.
+More precisely, for a given hypergeometric function, its Erdélyi-type integrals can generally express it as an integral of the product of two functions of the same type (see [33, p.
 2]).
@@
 Then in Section 3.2, we shall derive a curious integral representation for Appell function $F_{2}$ from Theorem 1.1.
-A further extension, which is related to Belitsky’s generalization of $F_{K}$ [4, p.
+A further extension, which is related to Belitsky’s generalization of $F_{K}$ [5, p.
 69], is considered in Section 3.3.
@@
 Let $f:\mathbb{C}^{k}\rightarrow\mathbb{C}$ be a function.
-The $k$-dimensional $q$-integral can be defined by (see, for example, [23, p.
-1] and [51, p.
+The $k$-dimensional $q$-integral can be defined by (see, for example, [24, p.
+1] and [59, p.
 241])
@@
 where $\mathbf{x}=(x_{1},\cdots,x_{k})$, $\sum_{\mathbf{n}\in\mathbb{Z}_{\geq 0}^{k}}\equiv\sum_{n_{1}=0}^{\infty}\cdots\sum_{n_{k}=0}^{\infty}$ and $\langle\mathbf{n}\rangle=n_{1}+\cdots+n_{k}$.
-The $q$-gamma function $\Gamma_{q}(x)$ is defined by [19, p. 20, Eq. (1.10.1)]
+The $q$-gamma function $\Gamma_{q}(x)$ is defined by [20, p. 20, Eq. (1.10.1)]
 $\Gamma_{q}(x)=\frac{(q;q)_{\infty}}{(q^{x};q)_{\infty}}(1-q)^{1-x}$
-and the $q$-beta function is defined by [19, p. 22, Eq. (1.10.13)]
+and the $q$-beta function is defined by [20, p. 22, Eq. (1.10.13)]
 $B_{q}(x,y)=\frac{\Gamma_{q}(x)\Gamma_{q}(y)}{\Gamma_{q}(x+y)}=(1-q)\frac{(q,q^{x+y};q)_{\infty}}{(q^{x},q^{y};q)_{\infty}}.$
-Note that $\lim_{q\rightarrow 1^{-}}\Gamma_{q}(x)=\Gamma(x)$ and $\lim_{q\rightarrow 1^{-}}B_{q}(x,y)=B(x,y)$. The $q$-integral representation for $B_{q}(x,y)$ is given by [19, p. 23, Eq. (1.11.7)]
+Note that $\lim_{q\rightarrow 1^{-}}\Gamma_{q}(x)=\Gamma(x)$ and $\lim_{q\rightarrow 1^{-}}B_{q}(x,y)=B(x,y)$. The $q$-integral representation for $B_{q}(x,y)$ is given by [20, p. 23, Eq. (1.11.7)]
 $B_{q}(x,y)=\int_{0}^{1}t^{x-1}\frac{(tq;q)_{\infty}}{(tq^{y};q)_{\infty}}\mathrm{d}_{q}t,\penalty 10000\ \penalty 10000\ \penalty 10000\ \Re(x)>0,\penalty 10000\ y\notin\mathbb{Z}_{\leq 0}.$
@@
 $\hskip 14.22636pt=\sum_{p=0}^{\infty}\frac{(\alpha_{2},\beta_{1};q)_{p}}{(\gamma_{3},q;q)_{p}}{}_{2}\phi_{1}\left[\begin{matrix}\beta_{1}q^{p},\alpha_{1}\\ \gamma_{1}\end{matrix};q,x\right]{}_{2}\phi_{1}\left[\begin{matrix}\alpha_{2}q^{p},\beta_{2}\\ \gamma_{2}\end{matrix};q,y\right]z^{p},$ (16)
-where $|x|<1$, $|y|<1$, $|z|<1$ and $_{r}\phi_{s}$ is the basic hypergeometric series defined by [19, p. 4, Eq. (1.2.22)]
+where $|x|<1$, $|y|<1$, $|z|<1$ and $_{r}\phi_{s}$ is the basic hypergeometric series defined by [20, p. 4, Eq. (1.2.22)]
 $_{r}\phi_{s}\left[\begin{matrix}a_{1},\cdots,a_{r}\\ b_{1},\cdots,b_{s}\end{matrix};q,z\right]:=\sum_{n=0}^{\infty}\frac{(a_{1},\cdots,a_{r};q)_{n}}{(b_{1},\cdots,b_{s},q;q)_{n}}\Big((-1)^{n}q^{\binom{n}{2}}\Big)^{1+s-r}z^{n}.$
@@
 $_{r}\widetilde{\phi}_{s}\left[\begin{matrix}a_{1},\cdots,a_{r}\\ b_{1},\cdots,b_{s}\end{matrix};q,z\right]={}_{r}\phi_{s}\left[\begin{matrix}q^{a_{1}},\cdots,q^{a_{r}}\\ q^{b_{1}},\cdots,q^{b_{s}}\end{matrix};q,z\right].$
-The only $q$-integral representation (of Bateman type) for $\Phi_{K}$ is given by Ernst [15, p. 10, Theorem 5]:
+The only $q$-integral representation (of Bateman type) for $\Phi_{K}$ is given by Ernst [16, p. 10, Theorem 5]:
 $\widetilde{\Phi}_{K}[\alpha_{1},\alpha_{2},\alpha_{2},\beta_{1},\beta_{2},\beta_{1};\gamma_{1},\gamma_{2},\gamma_{3};q,x,y,z]$
@@
 2 Preliminaries
-Gasper [18] discussed the $q$-analogues of Erdélyi’s integrals (5)–(7) and provided their discrete analogues. The integral (5) has the following $q$-analogue [18, p. 4, Eq. (1.13)]
+Gasper [19] discussed the $q$-analogues of Erdélyi’s integrals (7)–(9) and provided their discrete analogues. The integral (7) has the following $q$-analogue [19, p. 4, Eq. (1.13)]
 ${}_{2}\widetilde{\phi}_{1}\left[\begin{matrix}\alpha,\beta\\ \gamma\end{matrix};q,x\right]$ $=\int_{0}^{1}\frac{(xtq^{\alpha^{\prime}};q)_{\infty}}{(xt;q)_{\infty}}{}_{2}\widetilde{\phi}_{1}\left[\begin{matrix}\alpha-\alpha^{\prime},\beta\\ \lambda\end{matrix};q,xtq^{\alpha^{\prime}}\right]$
@@
 We have $\int_{0}^{1}\mathrm{d}\mu_{\eta-\lambda,\gamma-\lambda,\gamma-\lambda+\eta-\nu,\nu}(t;q)=1$ and $\lim_{q\rightarrow 1^{-}}\mathfrak{m}_{\alpha,\beta,\gamma,\eta}(t;q)=\mathfrak{m}_{\alpha,\beta,\gamma,\eta}(t)$.
-Then the $q$-analogue of (7) is given by [18, p.
+Then the $q$-analogue of (9) is given by [19, p.
 4, Eq.
@@
 $\cdot{}_{3}\phi_{2}\left[\begin{matrix}\mu/\lambda,\gamma/\lambda,q^{k-n}\\ \gamma\mu/(\lambda\nu),q^{1-n}/\lambda\end{matrix};q,q^{1-k}/\nu\right]{}_{4}\phi_{3}\left[\begin{matrix}\alpha,\beta,\mu,q^{-k}\\ \lambda,\nu,\delta\end{matrix};q,q\right].$ (24)
-Finally, we also need the following $q$-hypergeometric function of three variables (see [9, p. 96] and [40, p. 35]):
+Finally, we also need the following $q$-hypergeometric function of three variables (see [10, p. 96] and [41, p. 35]):
 $\phi^{(3)}[x,y,z]$ $\equiv\phi^{(3)}\left[\begin{matrix}(a)::(b);(b^{\prime});(b^{\prime\prime}):(c);(c^{\prime});(c^{\prime\prime})\\ (e)::(g);(g^{\prime});(g^{\prime\prime}):(h);(h^{\prime});(h^{\prime\prime})\end{matrix};q,x,y,z\right]$
@@
 $\hskip 42.67912pt\cdot F_{2}\left[a_{2},d_{1}-c_{1}+b_{1},d_{2};d_{1},d_{2};\frac{(1-v)y}{1-vy},\frac{wz}{1-vy}\right]\mathrm{d}\mu_{c_{1}-d_{1},d_{1}}(v)\mathrm{d}\mu_{b_{2},c_{2}-b_{2}}(w).$ (34)
-Then, by using [48, p. 306, Eq. (109)]
+Then, by using [49, p. 306, Eq. (109)]
 $F_{2}[a,b,b^{\prime};c,b^{\prime};y,z]=(1-z)^{-a}{}_{2}F_{1}\left[\begin{matrix}a,b\\ c\end{matrix};\frac{y}{1-z}\right].$ (35)
@@
 $\hskip 28.45274pt=(1-vy-wz)^{-a_{2}}{}_{2}F_{1}\left[\begin{matrix}a_{2},b_{1}+d_{1}-c_{1}\\ d_{1}\end{matrix};\frac{(1-v)y}{1-vy-wz}\right].$ (36)
-By using (33) and the Pfaff transformation [48, p. 300, Eq. (79)], we have
+By using (35) and the Pfaff transformation [49, p. 300, Eq. (79)], we have
 $F_{2}[a_{1}-a_{2},b_{1},b_{2}-d_{2};c_{1}-d_{1},b_{2}-d_{2};vy,wz]$
@@
 Remark 3.3.
-We can prove (3.2) directly by using the method described in Section 3.1. In addition, the integral (3.2) can also be derived from Manocha’s integral [34, p. 239, Eq. (13)]:
+We can prove (3.2) directly by using the method described in Section 3.1. In addition, the integral (3.2) can also be derived from Manocha’s integral [35, p. 239, Eq. (13)]:
 $F_{2}[a,b,c;d,e;y,z]=\int_{0}^{1}\int_{0}^{1}(1-vy-wz)^{-a^{\prime}}F_{2}[a-a^{\prime},b,c;\lambda,\eta;yv,zw]$
@@
 $F_{K}\left[\beta_{1},\alpha_{2},\alpha_{2},\alpha_{1},\beta_{2},\alpha_{1};\gamma_{1},\gamma_{2},\gamma_{3};x_{1},x_{2},x_{3}x_{4}\right].$
-Another significant example originates in physics. In 2018, Belitsky [4, p. 69] generalized Saran’s $F_{K}$-function of three variables to $L$ variables, i.e.,
+Another significant example originates in physics. In 2018, Belitsky [5, p. 69] generalized Saran’s $F_{K}$-function of three variables to $L$ variables, i.e.,
 $F_{K}\left[\begin{matrix}a_{1},b_{1},\cdots,b_{L-1},a_{2}\\ c_{1},\cdots,c_{L}\end{matrix};z_{1},\cdots,z_{L}\right]$
@@
 $|z_{1}|<1,\penalty 10000\ |z_{4}|<1\penalty 10000\ \penalty 10000\ \penalty 10000\ \text{and}\penalty 10000\ \penalty 10000\ \penalty 10000\ \frac{|z_{2}|}{1-|z_{1}|}+\frac{|z_{3}|}{1-|z_{4}|}<1.$
+Here it is worth mentioning that Khichi has introduced the multivariate hypergeometric functions $H_{B}^{(n)}$ ([49, p.
+308, Eq.
+(123)]), which appears quite similar to the function (3.3), although they are different in nature.
 By using the method from Section 3.1, we have the following theorem.
@@
 4.1 A Joshi-Vyas type theorem
-Inspired by the formula [21, p. 133] of Joshi and Vyas and its useful $q$-analogue [22, p. 646], we establish–using the measure in (19)–the following theorem of Joshi-Vyas type.
+Inspired by the formula [22, p. 133] of Joshi and Vyas and its useful $q$-analogue [23, p. 646], we establish–using the measure in (21)–the following theorem of Joshi-Vyas type.
 Theorem 4.1.
@@
 Remark 4.2.
-When $k=1$, the formula (4.1) reduces to the result obtained by Joshi and Vyas. It can also be considered as Ernst’s $q$-analogue for Exton’s result (see [14, p. 189, Theorem 3.1]).
+When $k=1$, the formula (4.1) reduces to the result obtained by Joshi and Vyas. It can also be considered as Ernst’s $q$-analogue for Exton’s result (see [15, p. 189, Theorem 3.1]).
 Letting $k=3$ and
@@
 Remark 4.4.
-This integral (4.3) is clearly a $q$-analogue of the integral by Luo et al. [32, Theorem 1]. In particular, when $x=0$, (4.3) reduces to
+This integral (4.3) is clearly a $q$-analogue of the integral by Luo et al. [33, Theorem 1]. In particular, when $x=0$, (4.3) reduces to
 $\widetilde{\Phi}_{2}[\alpha_{2},\beta_{2},\beta_{1};\gamma_{2},\gamma_{3};q,y,z]$ $=\int_{[0,1]^{2}}\widetilde{\Phi}_{0:2;2}^{1:2;2}\left[\begin{matrix}\alpha_{2}:\beta_{2},\eta_{2};\beta_{1},\eta_{3}\\ -:\nu_{2},\lambda_{2};\nu_{3},\lambda_{3}\end{matrix};q,yt_{2},zt_{3}\right]$
 $\hskip 42.67912pt\cdot\prod_{j=2}^{3}\mathrm{d}\mu_{\eta_{j}-\lambda_{j},\gamma_{j}-\lambda_{j},\gamma_{j}-\lambda_{j}+\eta_{j}-\nu_{j},\nu_{j}}(t_{j};q),$ (47)
-which is the $q$-analogue of Koschmieder’s integral representation for $F_{2}$ ([32, Corollary 1]).
-Here, $\Phi_{2}$ and $\Phi_{0:2;2}^{1:2;2}$ denote the $q$-Appell function [19, p.
+which is the $q$-analogue of Koschmieder’s integral representation for $F_{2}$ ([33, Corollary 1]).
+Here, $\Phi_{2}$ and $\Phi_{0:2;2}^{1:2;2}$ denote the $q$-Appell function [20, p.
 283, Eq.
-(10.2.6)] and the $q$-Kampé de Fériet function [48, p.
+(10.2.6)] and the $q$-Kampé de Fériet function [49, p.
 349], respectively.
@@
 Remark 4.6.
-This integral is the $q$-analogue of the integral by Luo and Raina [31, p. 14, Theorem 4.1].
+This integral is the $q$-analogue of the integral by Luo and Raina [32, p. 14, Theorem 4.1].
 In the next subsection, we shall show that Corollary 4.5 has a nice discrete analogue.
@@
 Let $f:\mathbb{C}\rightarrow\mathbb{C}$ be a function of one variable.
-We define the $q$-shift operator $\mathbb{E}_{q}$ by $\mathbb{E}_{q}f(x)=f(qx)$ $(0<q<1)$ [29, Definition 1.3].
+We define the $q$-shift operator $\mathbb{E}_{q}$ by $\mathbb{E}_{q}f(x)=f(qx)$ $(0<q<1)$ [30, Definition 1.3].
 Inductively, for any $n\in\mathbb{Z}_{\geq 1}$,
@@
 $\hskip 14.22636pt\cdot\sum_{p=0}^{k}\left[\begin{matrix}k\\ p\end{matrix}\right]_{q}(q^{\eta_{2}};q)_{k-p}(q^{\alpha_{2}-\eta_{2}};q)_{p}q^{(k-p)(\alpha_{2}-\eta_{2})},$
-where $\left[\begin{matrix}k\\ p\end{matrix}\right]_{q}$ denotes the $q$-binomial coefficient [19, p. 24]. Then the use of the $q$-binomial theorem [19, p. 25] gives
+where $\left[\begin{matrix}k\\ p\end{matrix}\right]_{q}$ denotes the $q$-binomial coefficient [20, p. 24]. Then the use of the $q$-binomial theorem [20, p. 25] gives
 $\mathcal{K}(q;x,y,z;u,v,w)$ $=\frac{(uxq^{\beta_{1}},vyq^{\alpha_{2}};q)_{\infty}}{(uxq^{\lambda_{3}},vyq^{\eta_{2}};q)_{\infty}}\sum_{k=0}^{\infty}\frac{(q^{\alpha_{2}};q)_{k}(wz)^{k}}{(uxq^{\beta_{1}},vyq^{\alpha_{2}},q;q)_{k}}$
@@
 Substituting (4.3) into (4.3) we obtain (4.10). ∎
+5 Conclusion
+In this paper, we continue the recent work of Luo, Xu, and Raina [33] on Erdélyi-type integrals for Saran’s function $F_{K}$, providing a new proof of one of their main results (Theorem 1.1).
+Building on this, we obtain an important generalization, which is applicable to the $L$-variable $F_{K}$ function (3.3) that arises in physics.
+Furthermore, some interesting $q$-analogues and discrete analogues are also discussed.
+By reviewing the literature on Erdélyi-type integrals of various hypergeometric functions over the past century, this paper compiles the table in Appendix 6, which reflects the overall landscape of research in this direction.
+As already clarified in Section 1, whether out of computational interest or due to their usefulness, it is highly valuable to continue exploring Erdélyi-type integrals for other hypergeometric functions not yet included in this table.
 6 Appendix
@@
 others Laine (1982) [29]
+Virchenko-Rumiantseva (2008) [55]
+Virchenko-Ovcharenko (2011) [54]
 $n$ variables Appell $F_{1}$ Manocha (1967) [35]
@@
 $q$-analogues the present paper
+others Volkodavov-Nikolaev (1993) [56]
+Author Contributions
+All the authors have contributed equally. All authors have read and approved the final manuscript.
 Funding
-No funding was received for conducting this study.
-Conflicts of interest
-The authors declare that there is no conflict of interest.
+The research of the second author is supported by National Natural Science Foundation of China (Grant No. 12001095).
 Data Availability
@@
 We are grateful to Peng-Cheng Hang for his careful reading of the initial manuscript, for correcting several typos, and for providing valuable suggestions.
+Conflicts of interest
+The authors declare that there is no conflict of interest.